### v2

# Pull-back of differential forms by multi-valued Sobolev maps, and the quasiregularity of the multi-valued inverse of a quasiregular map

Abstract.
We use Almgren’s framework of multi-valued maps to construct a multi-valued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\mathbb{R}^{n})$ of a quasiregular map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d$. We then develop a pull-back theory of differential forms on $\mathcal{A}_{d}(\mathbb{R}^{n})$ by Sobolev maps, and use it to show that the multi-valued inverse is a quasiregular $\omega$-curve (in the sense of Pankka) with respect to a natural $n$-form $\omega$ (suitably interpreted). The pull-back theory is of independent interest, and allows us to conclude e.g. higher Sobolev integrability and quasiminimality of the multi-valued inverse.
Key words and phrases:
Quasiregular maps, quasiregular curves, multi-valued maps, Almgren space, Sobolev differential forms, weak exterior derivative
2020 Mathematics Subject Classification:
30C65, 35R70, 53C23, 30L10
1. Introduction
1.1. Overview
A continuous map $f\in W^{1,n}_{loc}(M^{n},N^{k})$ between oriented Riemannian manifolds with $n\leq k$ is a $K$-quasiregular $\omega$-curve, if
$\|\omega\|\circ f\|Df\|^{n}\leq K\star f^{*}\omega\quad\textrm{ almost everywhere}.$ (1.1)
Here, $\omega$ is a (closed) $n$-form on $N$, $\star f^{*}\omega$ is the Hodge-star dual of $f^{*}\omega$, and $\|\omega\|$ is the pointwise comass of $\omega$. When $k=n$ and $\omega$ is the volume form on $N$, (1.1) recovers the notion of quasiregular (QR) maps, written more classically as
$\|Df\|^{n}\leq KJ_{f}\quad\textrm{ almost everywhere},$ (1.2)
where $J_{f}=\det(Df)$ and $Df$ is the weak differential of $f$. Quasiregular maps have a rich theory [26, 9], and enjoy many of the same analytic properties as quasiconformal (QC) maps. While they are not necessarily homeomorphisms, a seminal result of Reshetnyak [25] states that QR maps are open and discrete, and thus in particular locally injective outside a singular branch set of topological dimension at most n2n-2 [28]. We call an open discrete map a branched cover.
Quasiregular curves (QR curves) in turn were introduced by Pankka [24], and they generalize quasiregular maps by allowing the dimension of the target to be higher than that of the domain; for example pseudoholomorhic curves [5] are QR curves. Much of the classical theory of QR maps extends to QR curves [24, 21], with one notable exception: Reshetnyak’s theorem. Namely, QR curves need not be branched covers. Indeed, if $k>n$, a QR curve cannot be open, while the more non-trivial failure of discreteness follows from an example in [16].
Compared to QR maps, one of the challenges in QR curves is understanding how the choice of $\omega$ reflects the behaviour of $f$, see for example [7, 13] for some rigidity results and connections with calibrations.
In this paper, we identify a fruther connection between QR maps and QR curves: we construct a multi-valued inverse $\accentset{\leftarrow}{f}$ for a QR map $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d>0$, and show that $\accentset{\leftarrow}{f}$ is naturally a QR curve (see Theorem 1.2). In particular, QR maps have a local (multi-valued) inverse across the singular branch set (and not only outside it) with suitable Sobolev regularity. Earlier, related constructions include a push-forward operator [9, p. 263], [23, Section 4] and a generalized local inverse [26, Chapter II.5], [22, Section 8], both of which are important for obtaining modulus inequalities.
Our approach is based on Almgren’s space of unordered tuples [2] (also known as the symmetric product, cf. [6, Chapter 4.K]), and brings a new perspective to quasiregular theory; together with a pull-back theory by multi-valued maps developed in this paper (see Theorems 1.1 and 4.7) we obtain e.g. higher integrability of the generalized inverse [26, Chapter II.5] (see Corollary 1.3 and Theorem 5.3) as well as quasiminimality of the multi-valued inverse (see Theorem 5.4). We describe the construction and briefly discuss some its advantages over alternative approaches below.
1.2. Multi-valued inverse
Given a non-injective (proper) branched cover $f:\Omega\to\mathbb{R}^{n}$, there are several natural ways to interpret the set valued “inverse” map $f(\Omega)\ni y\mapsto f^{-1}(y)$. One is to consider the natural factorization $f=\bar{f}\circ P_{f}$, where $P_{f}:\Omega\to\widetilde{\Omega}_{f}:=\Omega/\sim$ is the projection onto the quotient space given by the equivalence relation $x\sim y$ if $f(x)=f(y)$, and $\bar{f}:\widetilde{\Omega}_{f}\to f(\Omega)$ the induced map, which in this case is a homeomorphism and thus invertible. Another approach is to view the set valued inverse as a map $f(\Omega)\to\mathcal{H}(\Omega)$ into the Hausdorff space of non-empty compact subsets of $\Omega$. A drawback of the first approach is that the target space of the inverse depends on $f$ and, more crucially, that $\bar{f}$ need not be geometrically quasiconformal unless $f$ is a BLD-map. The second approach, on the other hand, does not detect the multiplicities in the preimage, and $\mathcal{H}(\Omega)$ does not carry any natural differential forms, nor any bi-Lipschitz embedding in Euclidean space (cf. [2, Theorem 2.1]).
Instead, we give a construction based on Almgren’s space and local degree theory. Let $Q\geq 1$ be a natural number. The Almgren space $\mathcal{A}_{Q}(\mathbb{R}^{n})$ of unordered $Q$-tuples in $\mathbb{R}^{n}$ was used by Almgren in his seminal work on regularity of minimal currents [1], and was further explored by De Lellis et. al. [2]. It is a $nQ$-dimensional smooth orbifold, but not a manifold (unless $Q=1$), defined as
$\mathcal{A}_{Q}(\mathbb{R}^{n})=(\mathbb{R}^{n})^{Q}/S_{Q},\quad d_{\mathcal{A}}(\llbracket x\rrbracket,\llbracket y\rrbracket)=\inf_{\sigma\in S_{Q}}\Big(\sum_{j}^{d}|x_{j}-y_{\sigma(j)}|^{2}\Big)^{1/2},$
where $S_{Q}$ is the permutation group on $Q$ elements acting naturally on $(\mathbb{R}^{n})^{Q}$. (See Section 2.2 for more details.)
The link between the multi-valued inverse and Almgren space is the following simple observation based on local degree theory. Given a proper branched cover $f:\Omega\to\mathbb{R}^{n}$ of degree $d>0$ from an open set $\Omega\subset\mathbb{R}^{n}$, the local index $\iota(f,x)$ is defined for all $x\in\Omega$ and satisfies $\sum_{x\in f^{-1}(y)}\iota(f,x)=d$, $y\in f(\Omega)$. Thus, $f^{-1}(y)$ can be regarded as an unordered $d$-tuple in $\Omega$ (counted with multiplicities) for each $y\in f(\Omega)$. The multi-valued inverse $\accentset{\leftarrow}{f}$ of the proper branched cover $f:\Omega\to\mathbb{R}^{n}$ of finite degree $d>0$ is the map
$\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\mathbb{R}^{n}),\quad\accentset{\leftarrow}{f}(y):=\sum_{x\in f^{-1}(y)}\iota(f,x)\llbracket x\rrbracket.$ (1.3)
We refer the reader to Section 2.1 (see in particular Definition 2.1) for the notation and further details.
1.3. Statement of results
While it is not a manifold, $\mathcal{A}_{Q}(\mathbb{R}^{n})$ nevertheless carries a natural $n$-form given by the trace of the volume form of $\mathbb{R}^{n}$; that is, the $n$-form
$\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}}):=\sum_{j}^{Q}P_{j}^{*}(\operatorname{vol}_{\mathbb{R}^{n}})$ (1.4)
on $(\mathbb{R}^{n})^{Q}$ (where $P_{j}$ is the projection to the $j$th factor) is $S_{Q}$-invariant: $\sigma^{*}\omega=\omega$ for all $\sigma\in S_{Q}$. Using $S_{Q}$-invariance, we develop a pull-back theory of multi-valued maps (see Sections 4.1 and 4.2), which is of independent interest.
Theorem 1.1.
Let $n,m,Q\geq 1$ be natural numbers, $U\subset\mathbb{R}^{m}$ open, and $f\in W_{loc}^{1,p}(f(\Omega),\mathcal{A}_{Q}(\mathbb{R}^{n}))$. If $\omega$ is a $S_{Q}$-invariant smooth $k$-form on $(\mathbb{R}^{n})^{Q}$, and either $k+1\leq p$ of $\mathrm{d}\omega=0$ and $k\leq p$, then $\omega$ has a well-defined pull-back $f^{*}\omega\in L^{p/k}_{loc}(U,\bigwedge^{k}\mathbb{R}^{m})$ and moreover
$\mathrm{d}(f^{*}\omega)=f^{*}(\mathrm{d}\omega)$
weakly.
In particular, if $f:U\to\mathcal{A}_{Q}(\mathbb{R}^{n})$ is a multi-valued (locally) Lipschitz map, the pull-back $f^{*}\omega$ of any $S_{Q}$-invariant smooth $k$-form $\omega$ is a (locally) flat $k$-form on $U$ in the sense of Whitney (cf. Theorem 4.7).
Applying the definition of pull-back in the QR context we show that, if $f:\Omega\to\mathbb{R}^{n}$ is a quasiregular map, the multi-valued inverse is a quasiregular curve.
Theorem 1.2.
Let $f:\Omega\to\mathbb{R}^{n}$ be a proper quasiregular map of finite degree $d$. Then the multivalued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\Omega)$ is an $\omega$-quasiregular curve, where $\omega=\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}})$ is given by (1.4). More precisely, $\accentset{\leftarrow}{f}\in W_{loc}^{1,n}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$, and
$\|\omega\|\circ\accentset{\leftarrow}{f}\ |D\accentset{\leftarrow}{f}|^{n}\leq d^{n/2-1}K_{I}(f)\star\accentset{\leftarrow}{f}^{*}\omega\quad\textrm{almost everywhere on }\Omega.$
Here $K_{I}(f)$ is the inner distortion of the quasiregular map $f$, see Section 2.1. Theorems 1.1 and 1.2 together with standard arguments found e.g. in [21] yield in particular the higher integrability of the generalized inverse of [26, Chapter II.5]. Recall that, given a proper quasiregular map $f:\Omega\to\mathbb{R}^{n}$, the generalized inverse is the map $g_{f}:f(\Omega)\to\mathbb{R}^{n}$ given by
$g:f(\Omega)\to\mathbb{R}^{n},\quad g(y):=\sum_{x\in f^{-1}(y)}\iota(f,x)x.$ (1.5)
We record this higher integrability result in Corollary 1.3 below, along with the somewhat surprising fact that the generalized inverse (1.5) is also a quasiminimizer of the Dirichlet energy. While both statements follow from the corresponding results for the multi-valued inverse (Theorems 5.3 and 5.4), the latter requires further analysis of the structure of Almgren’s space, presented in Appendix A. Indeed in Appendix A we show that $\mathcal{A}_{d}(\mathbb{R}^{n})$ splits off a factor $\mathbb{R}^{n}$, and cannot split off any higher dimensional Euclidean factors, cf. Proposition A.1.
Corollary 1.3.
Let $f:\Omega\to\mathbb{R}^{n}$ be a proper quasiregular map, and $g_{f}$ its generalized inverse. Then
  • (i)
    there exists $p>n$ such that $g_{f}\in W^{1,p}_{loc}(f(\Omega),\mathbb{R}^{n})$;
  • (ii)
    if $h\in W^{1,n}_{loc}(f(\Omega),\mathbb{R}^{n})$ is such that $\{h\neq g_{f}\}$ is essentially contained in a compact set $K\subset f(\Omega)$, then
    $\int_{K}\|Dg_{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh\|^{n}\mathrm{d}x.$
The multi-valued inverse can also be seen as a homeomorphism $\accentset{\leftarrow}{f}:f(\Omega)\to\Omega_{f}$ onto its image $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega))\subset\mathcal{A}_{d}(\mathbb{R}^{n})$ (Lemma 3.1). Thus, although $\mathcal{A}_{d}(\mathbb{R}^{n})$ is not a manifold, $\Omega_{f}$ is a metric $n$-manifold when equipped with the metric from $\mathcal{A}_{d}(\mathbb{R}^{n})$. In the following theorem we obtain that $\accentset{\leftarrow}{f}:f(\Omega)\to\Omega_{f}$ is geometrically quasiconformal.
Theorem 1.4.
Let $f:\Omega\to\mathbb{R}^{n}$ be a proper quasiregular map of finite degree $d$, and equip $\Omega_{f}=\accentset{\leftarrow}{f}(f(\Omega))$ with the metric from $\mathcal{A}_{d}(\mathbb{R}^{n})$ and the Hausdorff $n$-measure on $\Omega_{f}$.
  • (1)
    $\Omega_{f}$ is $n$-rectifiable, upper Ahlfors $n$-regular and satisfies the infinitesimal $n$-Poincaré inequality;
  • (2)
    $\accentset{\leftarrow}{f}:f(\Omega)\to\Omega_{f}$ is geometrically quasiconformal. More precisely,
    $\frac{1}{K_{I}K_{O}}\operatorname{Mod}_{n}\Gamma\leq\operatorname{Mod}_{n}\accentset{\leftarrow}{f}(\Gamma)\leq K_{I}K_{O}\operatorname{Mod}_{n}\Gamma$
    fir any path family $\Gamma$ in $f(\Omega)$.
We refer to (6.1) in Section 6.1 for the definition of infinitesimal Poincaré inequality and further commentary. We mention here that the infinitesimal Poincaré inequality is not a quantitative condition.
1.4. Main ideas and further discussion
Theorem 1.1 follows by an approximation argument from its Lipschitz counterpart, Theorem 4.7, whose proof uses induction, decomposition of multi-valued maps (cf. [2, Proposition 1.6]), and a somewhat delicate approximation scheme. We outline the main idea.
In order to apply induction in the proof of Theorem 4.7, we will need to approximate a given multi-valued Lipschitz map by maps which are locally decomposable. A given map $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ fails to be locally decomposable on the set $F:=f^{-1}(\operatorname{diag}\mathcal{A}_{d}(\mathbb{R}^{n}))$. The restriction of $f$ to $F$ agrees with the map $d\llbracket b(f)\rrbracket$, where $b:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathbb{R}^{n}$ is the barycenter map
$b(\llbracket x_{1},\ldots,x_{d}\rrbracket)=\frac{x_{1}+\cdots+x_{d}}{d},$ (1.6)
which is $\frac{1}{\sqrt{d}}$-Lipschitz, see Corollary A.2. The equality $f|_{F}=d\llbracket b(f)\rrbracket|_{F}$ implies that the differentials of the two functions agree almost everywhere on $F$. However, since $F$ is closed and not open, we need an approximation argument for the induction step to work. Our approximation scheme consists of a careful interpolation between $f$ and $d\llbracket b(f)\rrbracket$ in a neighbourhood of $F$.
The main step in the proof of Theorem 1.2 are the Sobolev estimates, which we obtain in Theorem 3.7. Our argument here uses Vitali’s covering theorem, which is valid in high generality, and thus Theorem 3.7 probably holds e.g. in the setting of [22].
The proof of Theorem 1.4 employs a similar induction argument as in Theorem 4.7 together with some metric quasiconformal theory. We remark here that a large part of metric quasiconformal theory is based on the Ahlfors regularity of the spaces under consideration. Indeed, in our setting this property would imply the Loewner property of the image set $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega)$ by a (deep) result of Semmes [27], which in turn implies the equivalence of the various notions of quasiconformality in the metric setting [11]. In this paper, we establish the upper Ahlfors $n$-regularity of $\Omega_{f}$ (Proposition 6.2), and an infinitesimal Poincaré inequality. It would be interesting to analyze the geometry of $\Omega_{f}$ further, in particular whether $\Omega_{f}$ admits local bi-Lipschitz parametrizations. By Almgren’s bi-Lipschitz embedding $\mathcal{A}_{d}(\mathbb{R}^{n})\hookrightarrow\mathbb{R}^{N}$, $\Omega_{f}$ can be regarded as a subset of Euclidean space, and the existence of bi-Lipschitz parametrizations is linked with the existence of Cartan–Whitney presentations in a suitable Sobolev class [8, [Theorem 1.2].
Acknowledgements
The research in this manuscript was supported by Research Council of Finland grant no. 355122. The author thanks Pekka Pankka, Ilmari Kangasniemi and Toni Ikonen for useful discussions on the manuscript.
2. Preliminaries
2.1. Branched covers and quasiregular maps
For the notion of local degree and other topological preliminaries we refer the reader to [26, Chapter I.4]. In this paper we call a continuous open discrete map a branched cover. A precompact open set $U\subset X$ is a normal neighbourhood of $f$ if $\partial f(U)=f(\partial U)$. A branched cover $f:X\to Y$ between topological $n$-manifolds has a well-defined local index $\iota(f,x)$ for every $x\in X$, satisfying
$\deg(f,U,y)=\sum_{x\in f^{-1}(y)\cap U}\iota(f,x).$
More generally, for any precompact open $U\subset X$, the map
$f(U)\setminus f(\partial U)\to\mathbb{Z},\quad y\mapsto\sum_{x\in f^{-1}(y)}\iota(f,x)$
is constant on each connected component of $f(U)\setminus f(\partial U)$. Moreover, $f$ is a local homeomorphism in some neighbourhood of a given point $x\in X$ if and only if $\iota(f,x)=\pm 1$. The branch set
$B_{f}=\{x\in X:|\iota(f,x)|>1\}$
is a closed set of topological dimension at most n2n-2 [28]. It follows that branched covers are either sense preserving ($\iota(f,\cdot)\geq 1$ everywhere) or sense reversing ($\iota(f,\cdot)\leq-1$ everywhere). If $f:X\to Y$ is a proper branched cover of degree finite $d>0$, then
$\sum_{x\in f^{-1}(y)}\iota(f,x)=d\quad\textrm{for all }y\in f(X).$ (2.1)
Definition 2.1.
Let $f:X\to Y$ be a proper branched cover of finite degree $d>0$ between topological $n$-manifolds. The multi-valued inverse $\accentset{\leftarrow}{f}:f(X)\to\mathcal{A}_{d}(X)$ of $f$ is defined by
$\accentset{\leftarrow}{f}(y)=\sum_{x\in f^{-1}(y)}\iota(f,x)\llbracket x\rrbracket$
It follows from (2.1) that $\accentset{\leftarrow}{f}$ is well-defined.
Recall that a Sobolev map $f\in W^{1,n}_{loc}(\Omega,\mathbb{R}^{n})$ ($\Omega\subset\mathbb{R}^{n}$ open, $n\geq 2$) is called quasiregular, if there exists $K$ such that (1.2) holds. See [26] for an excellent introduction to quasiregular maps. The least constant $K$ in (1.2) is called the outer distortion and denoted $K_{O}(f)$. The inner distortion $K_{I}(f)$ is the smallest constant $K$ such that
$J_{f}\leq K\min_{v\in S^{n-1}}\{\|Df(v)\|^{n}\}\ \textrm{ almost everywhere}.$
If $f$ is quasiconformal, i.e. a quasiregular homeomorphism, then the inverse $f^{-1}$ is quasiconformal and $K_{O}(f^{-1})=K_{I}(f)$, $K_{I}(f^{-1})=K_{O}(f)$. We will make use of this fact without further mention in the sequel.
By the seminal result of Reshetnyak [25], non-constant quasiregular maps are continuous, open and discrete. In particular they are sense-preserving and $J_{f}\in(0,\infty)$ almost everywhere.
2.2. Space of unordered tuples
Let $X$ be a metric space and $d\geq 1$ a natural number. The Almgren space – or symmetric product – $\mathcal{A}_{d}(X)$ is the space $X^{d}/S_{d}$ equipped with the metric
$d_{\mathcal{A}}(\bar{x},\bar{y})=\min_{\sigma\in S_{d}}\Big(\sum_{j}^{d}d(x_{j},y_{\sigma(j)})^{2}\Big)^{1/2},$ (2.2)
for $\bar{x}=\llbracket x_{1},\ldots,x_{d}\rrbracket,\ \bar{y}=\llbracket y_{1},\ldots,y_{d}\rrbracket$. Here, the action by isometries on $X^{d}$ of the symmetric group $S_{d}$ is given by $\sigma\cdot(x_{1},\ldots,x_{d})=(x_{\sigma^{-1}(1)},\ldots,x_{\sigma^{-1}(d)})$. The space $\mathcal{A}_{d}(X)$ can alternatively be described as a subset of the Wasserstein space $\mathcal{P}_{2}(X)$, consisting of probability measures $\mu\in\mathcal{P}_{2}(X)$ such that $\operatorname{spt}(\mu)\leq d$ and $d\cdot\mu(\{x\})\in\mathbb{N}$ for all $x\in X$ (note however that the metric inherited from $\mathcal{P}_{2}(X)$ differs from (2.2) by a factor of $1/\sqrt{d}$). Following [2], we denote elements of $\mathcal{A}_{d}(X)$ either by $\llbracket x_{1},\ldots,x_{d}\rrbracket$ or by $\sum_{j}^{m}a_{j}\llbracket x_{j}\rrbracket$, where the numbers $a_{j}>0$ satisfy $\sum_{j}^{m}a_{j}=d$.
Multi-valued differentiability
The following Rademacher’s theorem for multi-valued Lipschitz maps is from [2, Theorem 1.13]. In the statement $m,n,d\geq 1$ are natural numbers and $U\subset\mathbb{R}^{m}$ an open set.
Theorem 2.2.
Let $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ be locally Lipschitz. Then $f$ is differentiable for almost every $x_{0}\in U$ in the following sense: there are linear maps $L_{1},\ldots,L_{d}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ so that
$T_{x_{0}}f(x):=\llbracket f_{1}(x_{0})+L_{1}(x-x_{0}),\ldots,f_{d}(x_{0})+L_{d}(x-x_{0})\rrbracket$
satisfies $T_{x_{0}}f(x_{0})=\llbracket f_{1}(x_{0}),\ldots,f_{d}(x_{0})\rrbracket=f(x_{0})$ and
  • (i)
    $\lim_{x\to x_{0}}\frac{d_{\mathcal{A}_{d}}(f(x),T_{x_{0}}f(x))}{|x-x_{0}|}=0$, and
  • (ii)
    $L_{i}=L_{j}$ if $f_{i}(x_{0})=f_{j}(x_{0})$.
The differential $Df:=\llbracket L_{1},\ldots,L_{d}\rrbracket$ can be regarded as an element in $\mathcal{A}_{d}(\mathbb{R}^{n\times n})$, but note that this does not determine it unambiguously, cf. [2, Remark 1.11]. If $f_{i}:U\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$, $i=0,1$, are differentiable at $x\in U$ and $d_{0}+d_{1}=d$, then the map $f:=\llbracket f_{0},f_{1}\rrbracket:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is differentiable at $x\in U$, and $D_{x}f=\llbracket D_{x}f_{0},D_{x}f_{1}\rrbracket$.
Multi-valued Sobolev maps
We refer the reader to [2, Chapter 2.2 and Chapter 4] for the theory of Sobolev maps $W^{1,p}_{loc}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$ and mention here that Sobolev maps are approximately differentiable almost everywhere [2, Corollary 2.7]. This follows from the following Lusin-type approximations by Lipschitz maps [2, Proposition 2.5].
Proposition 2.3.
Let $f\in W^{1,p}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$. For every $\lambda>0$ there exists a $\lambda$-Lipschitz map $f_{\lambda}:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ such that
$\lambda^{p}\mathcal{L}^{m}(\{f\neq f_{\lambda}\})\leq C\int_{\{f\neq f_{\lambda}\}}(d_{\mathcal{A}}(f,d\llbracket 0\rrbracket)^{p}+|Df|^{p})\mathrm{d}x.$ (2.3)
Proof.
The claim is stated with $\{f\neq f_{\lambda}\}$ replaced by $U$ on the right hand side of (2.3) in [2, Proposition 2.5], but the proof and references therein yield the inequality stated here. ∎
Metric Jacobian
Let $s$ be a seminorm on $\mathbb{R}^{n}$. The Jacobian $\mathbf{J}(s)$ is the unique number satisfying $\mathcal{H}^{n}_{s}=\mathbf{J}(s)\mathcal{L}^{n}$, where $\mathcal{H}^{n}_{s}$ is zero if $s$ is degenerate, and the Hausdorff $n$-measure with respect to the metric induced by $s$ otherwise. If $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is (approximately) differentiable at $x\in U$, we denote
$\mathbf{J}f(x)=\mathbf{J}(\operatorname{md}_{x}f),$
where $\operatorname{md}_{x}f$ is the seminorm on $\mathbb{R}^{n}$ given by $\operatorname{md}_{x}f(v)=|D_{x}f(v)|$ for $v\in\mathbb{R}^{n}$. [^1] Note that this is well-defined since the norm of $D_{x}f(v)$ is independent of the permutation of order of the elements in $D_{x}f=(D_{x}f_{1},\ldots,D_{x}f_{d})$. We refer to [18] where the metric differential and Jacobian were introduced in order to prove an area formula for maps into metric spaces (see also [17]). We record the following identity for future use.
Lemma 2.4.
Suppose $L=\llbracket L_{1},\ldots,L_{d}\rrbracket:\mathbb{R}^{n}\to\mathcal{A}_{d}(\mathbb{R}^{n})$, where $L_{1},\ldots,L_{d}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ are linear maps, and let $s$ be the seminorm given by $s(v)=\Big(\sum_{j}^{d}|L_{j}(v)|^{2}\Big)^{1/2}$ for $v\in\mathbb{R}^{n}$. Then
$\mathbf{J}(s)=\det(L_{1}^{T}L_{1}+\cdots+L_{d}^{T}L_{d})^{1/2}.$
In particular, if $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is differentiable a.e., then
$\mathbf{J}f=\det\Big(\sum_{j}^{d}Df_{j}^{T}Df_{j}\Big).$
Proof.
If $f$ is differentiable almost everywhere, then it is metrically differentiable almost everywhere and
$\operatorname{\operatorname{md}}_{x}f(v)=\Big(\sum_{j}^{d}|Df_{j}(v)|^{2}\Big)^{1/2},\quad v\in\mathbb{R}^{m}$
for a.e. $x\in U$. Thus the second claim readily follows from the first.
To prove the first, it suffices to consider the linear map $L=(L_{1},\ldots,L_{d}):\mathbb{R}^{n}\to(\mathbb{R}^{n})^{d}$. By the area formula on Euclidean spaces, we have that $J_{n}(L)L_{\ast}\mathcal{L}^{n}=\mathcal{H}^{n}|_{\operatorname{Im}(L)}$, where
$J_{n}(L)=|\det(L^{T}L)|^{1/2}=|\det(L_{1}^{T}L_{1}+\cdots+L_{d}^{T}L_{d})|^{1/2}$
Now $J_{n}(L)\neq 0$ if and only if $L$ has full rank, in which case $s$ is a genuine norm. Note that $L:(\mathbb{R}^{n},s)\to(\operatorname{Im}(L),\|\cdot\|_{Eucl})$ is an isometry, and thus $L_{\ast}\mathcal{H}^{n}_{s}=\mathcal{H}^{n}|_{\operatorname{Im}(L)}$. Since $L:\mathbb{R}^{n}\to\operatorname{Im}(L)$ is bijective, this yields $\mathcal{H}^{n}_{s}=J_{n}(L)\mathcal{L}^{n}$. ∎
3. Multivalued inverse
3.1. Basic properties for branched covers
Throughout this subsection, $f:X\to Y$ is a proper branched cover of degree $d$ between metric spaces $X,Y$ homeomorphic to topological $n$-manifolds, $n\geq 2$. We start with a simple lemma.
Lemma 3.1.
The generalized inverse $\accentset{\leftarrow}{f}:f(X)\to\mathcal{A}_{d}(X)$ is a homeomorphism onto its image.
In the proof we use the following notation. Let $U_{f}(x,r)$ be the connected component of $f^{-1}B(f(x),r)$ containing $x$. For each $y\in f(X)$, there exists $r_{0}$ such that $U_{f}(x,r)$ is a normal neighbourhood of $x$ for all $x\in f^{-1}(y)$ and $r\leq r_{0}$. In particular the closure of $U_{f}(x,r)$ is compact and $f(\partial U_{f}(x,r))=\partial B(f(x),r)$.
Proof.
It is clear that $\accentset{\leftarrow}{f}$ is injective. To prove continuity, observe that
$\lim_{r\to 0}\operatorname{diam}U_{f}(x,r)=0$
for all $x\in X$. Indeed,
$\bigcap_{j}\overline{U}_{f}(x,2^{-j})=\overline{U}_{f}(x,r_{0})\cap f^{-1}\Big(\bigcap_{j}\overline{B}(f(x),2^{-j})\Big)=\{x\},$
and since the sets $\overline{U}_{f}(x,2^{-j})$ are compact and nested for large enough $j$, it follows that the diameters must tend to zero. (See also [26, Chapter I, Lemma 4.9].)
Fix $y\in f(X)$ and $\varepsilon>0$. Let $\delta>0$ be small enough such that $\operatorname{diam}U_{f}(x,\delta)<\varepsilon$ for all $x\in f^{-1}(y)$ (note that $f^{-1}(y)\leq d$). For $y^{\prime}\in B(y,\delta)$ we may group the preimages and obtain
$f^{-1}(y^{\prime})=\sum_{x\in f^{-1}(y)}\sum_{x^{\prime}\in f^{-1}(y^{\prime})\cap U_{x}}\iota(f,x^{\prime})\delta_{x^{\prime}}$
(where $U_{x}:=U_{f}(x,\delta)$). Moreover $\iota(f,x)=\sum_{x^{\prime}\in f^{-1}(y^{\prime})\cap U_{x}}\iota(f,x^{\prime})$. Consequently
$d_{\mathcal{A}_{d}}(f^{-1}(y),f^{-1}(y^{\prime}))^{2}$ $\leq\sum_{x\in f^{-1}(y)}\sum_{x^{\prime}\in f^{-1}(y^{\prime})\cap U_{x}}\iota(f,x^{\prime})d(x,x^{\prime})^{2}$
$\leq\sum_{x\in f^{-1}(y)}\iota(f,x)\varepsilon^{2}=d\varepsilon^{2}.$
This proves the claimed continuity. We complete the proof by showing $\accentset{\leftarrow}{f}^{-1}:\accentset{\leftarrow}{f}(f(X))\to f(X)$ is continuous. Suppose $z_{k}=\accentset{\leftarrow}{f}(y_{k})\to z=\accentset{\leftarrow}{f}(y)$. For small $r>0$ we have $f^{-1}B(y,r)=\bigcup_{x\in f^{-1}(y)}U_{f}(x,r)$ where each $U_{f}(x,r)=:U_{x}$ is a normal neighbourhood of $x$. For each $k$ write $f^{-1}(y_{k})=\bigcup_{x\in f^{-1}(y)}(f^{-1}(y_{k})\cap U_{x})$. Since $\accentset{\leftarrow}{f}(y_{n})\to\accentset{\leftarrow}{f}(y)$, for large $k$ we have that $f^{-1}(y_{k})\cap U_{x}\neq\varnothing$ for all $x\in f^{-1}(y)$ and
$d_{\mathcal{A}}(z_{k},z)^{2}=\sum_{x\in f^{-1}(y)}\sum_{x_{k}\in f^{-1}(y_{k})\cap U_{x}}d(x_{k},x)^{2}\stackrel{{\scriptstyle k\to\infty}}{{\longrightarrow}}0.$
From this it follows that for each $x\in f^{-1}(y)$ there exists $x_{k}\in f^{-1}(y_{k})$ with $x_{k}\to x$, which in turn implies $y_{k}=f(x_{k})\to f(x)=y$ by the continuity of $f$. This proves that $\accentset{\leftarrow}{f}^{-1}$ is continuous. ∎
The next lemma relates the singular set of $\mathcal{A}_{d}(X)$ and the branch set of $f$. Here, the singular set $\operatorname{sing}\mathcal{A}_{d}(X)$ is the union of the sets
$\operatorname{sing}_{k}\mathcal{A}_{d}(X):=\{\llbracket x_{1},\ldots,x_{d}\rrbracket:\ x_{1}=\cdots=x_{k}\},\quad k=2,\ldots,d.$
We also denote $\operatorname{sing}_{d}\mathcal{A}_{d}(X)=\operatorname{diag}\mathcal{A}_{d}(X)$.
Lemma 3.2.
We have that
$\accentset{\leftarrow}{f}^{-1}(\operatorname{sing}(\mathcal{A}_{d}(X)))=f(B_{f}).$
Proof.
A point $y\in f(X)$ is in $f(B_{f})$ if and only if $f^{-1}(y)\cap B_{f}\neq\varnothing$. On the other hand, $F(y)\in\operatorname{sing}(\mathcal{A}_{d}(X))$ if and only if $\iota(f,x)>1$ for some $x\in f^{-1}(y)$. These two conditions are equivalent. ∎
The following proposition will be useful later on. It is based on a theorem of Floyd [4], see also [19, Proposition 2.6], which states that compact paths admit lifts by a proper branched cover.
Proposition 3.3.
Let $\gamma:[0,1]\to f(X)$ be a continuous path. Then
$\accentset{\leftarrow}{f}\circ\gamma=\sum_{x\in f^{-1}(\gamma_{0})}\sum_{j}^{\iota(f,x)}\llbracket\gamma_{x}^{j}\rrbracket,$ (3.1)
where $\gamma_{x}^{j}$ ($j=1,\ldots,\iota(f,x)$) are the lifts of $\gamma$ by $f$ starting at $x$, i.e. $f\circ\gamma_{x}^{j}=\gamma$, $\gamma_{x}^{j}(0)=x$.
Proof.
By arguing as in [26, Chapter II, Corollary 3.4] using [4] (cf. [19, Proposition 2.6]), we see that $\gamma$ has $d$ total $f$-lifts $\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}$ which satisfy $\{j:\ \tilde{\gamma}_{j}(t)=x\}=\iota(f,x)$ for all $x\in f^{-1}(\gamma_{t})$. Relabeling these as $\gamma_{x}^{j}$ for $j=1,\ldots,\iota(f,x)$ and each $x\in f^{-1}(\gamma_{0})$, this implies $\accentset{\leftarrow}{f}(\gamma_{t})=\sum_{x\in f^{-1}(y)}\sum_{j}^{\iota(f,x)}\llbracket\gamma_{x}^{j}(t)\rrbracket$ for all $t\in[0,1]$, proving (3.1). ∎
Proposition 3.4.
Suppose $s\geq 0$ and $E\subset\accentset{\leftarrow}{f}(f(X))$. Then
$\mathcal{H}^{s}((\accentset{\leftarrow}{f}\circ f)^{-1}E)\leq d\cdot\mathcal{H}^{s}(E).$
Proof.
Denote $F=\accentset{\leftarrow}{f}\circ f:X\to\mathcal{A}_{d}(X)$ and $X_{f}=F(X)$. Note that $F:X\to X_{f}$ is a branched cover of degree $d$. First suppose $E$ is contained in a compact set $K\subset X_{f}$. For each $z_{0}\in K$ let $r_{0}$ be so small that
$F^{-1}B(z_{0},r_{0})=\bigcup_{x_{0}\in F^{-1}(z_{0})}U_{F}(x_{0},r_{0})$
where $U_{F}(x_{0},r_{0})$ is a normal neighbourhood of $x_{0}$ for each $x_{0}\in F^{-1}(z_{0})$. Suppose $\{E_{i}\}$ is a $\delta$-cover of $E$, where $\delta$ is smaller than half the Lebesgue number of the cover $\{B(z_{0},r_{0})\}_{z_{0}\in K}$. Then $E_{i}\subset B(z_{i},r_{i})$ for some $z_{i}\in K$ with $m=m_{i}$ preimages $x_{1},\ldots,x_{m}$ under $F$ ($m\leq d$). Thus
$F^{-1}(E_{i})=\bigcup_{k}^{m}E_{i,k},\quad E_{i,k}=F^{-1}(E_{i})\cap U_{F}(x_{k},r_{i}).$
Moreover, since $d_{\mathcal{A}}(F(x),F(y))\geq d(x,y)$ for $x,y\in U_{F}(x_{k},r_{i})$, it follows that $\operatorname{diam}E_{i,k}\leq\operatorname{diam}E_{i}$ for all $i$ and $k=1,\ldots,m_{i}$. The two observations above imply that $\{E_{i,k}\}$ is a $\delta$-cover of $F^{-1}E$, yielding
$\mathcal{H}_{\delta}^{s}(F^{-1}E)\leq\sum_{i}\sum_{k=1}^{m_{i}}(\operatorname{diam}E_{i,k})^{s}\leq\sum_{i}d\cdot(\operatorname{diam}E_{i})^{s}.$
Taking infimum over all $\delta$-covers and sending $\delta$ to zero we obtain
$\mathcal{H}^{s}(F^{-1}E)\leq d\mathcal{H}^{s}(E).$
Exhausting $X_{f}$ by compact sets and using basic convergence theorems of measures we the claim follows for general $E\subset X_{f}$. ∎
Push-forward
We define a push-forward operator $f_{\ast}$ as follows: given a Borel function $g:X\to[-\infty,\infty]$, set
$f_{\ast}g(y)=\sum_{x\in f^{-1}(y)}\iota(f,x)g(x),\quad y\in f(X).$ (3.2)
We record here the fact that the push-forward preserves continuity and measurability. This should be compared to [9, Lemma 14.31] and [23, Lemma 4.1] where it is shown that it preserves Sobolev regularity (resp. Lipschitz regularity when $f$ is BLD). Later on the push-forward will prove useful in connection with area formulae, see Lemma 3.6 and Proposition 3.9.
Lemma 3.5.
If $g\in C(X)$, then $f_{\ast}g\in C(f(X))$. If $g:X\to\mathbb{R}\cup\{\infty\}$ is Borel, then $f_{\ast}g:f(X)\to\mathbb{R}\cup\{\infty\}$ is Borel.
Proof.
Given a function $g$ on $X$, define $g_{\mathcal{A}_{d}}$ on $\mathcal{A}_{d}(X)$ by
$g_{\mathcal{A}_{d}}(\llbracket x_{1},\ldots,x_{d}\rrbracket)=g(x_{1})+\cdots+g(x_{d})$
Observe that $g_{\mathcal{A}_{d}}$ is continuous (resp. Borel) if $g$ is continuous (resp. Borel). Finally note that $f_{\ast}g=g_{\mathcal{A}_{d}}\circ\accentset{\leftarrow}{f}$, from which both claims follow. ∎
3.2. Sobolev estimates
In this subsection we specialize to quasiregular maps on Euclidean domains. Throughout this subsection we assume that $f:\Omega\to\mathbb{R}^{n}$ is a proper quasiregular map of degree $d$, where $\Omega\subset\mathbb{R}^{n}$ is a domain and $n\geq 2$. The following lemma is a reformulation of the area formula using the push-forward operator. Together with Theorem 3.7 below it yields the local Sobolev regularity of the multi-valued inverse.
Lemma 3.6.
If $g:X\to[0,\infty]$ is Borel, then
$\int_{\Omega^{\prime}}f_{\ast}g(y)\mathrm{d}\mathcal{H}^{n}(y)=\int_{\Omega}g\mathbf{J}f\mathrm{d}\mathcal{H}^{n}.$
Proof.
Since $|B_{f}|=0$, we have that $\iota(f,x)=1$ $\mathcal{H}^{n}$-a.e. $x\in\Omega$. Recall that a quasiregular map has the property (N) and thus the co-area formula holds for $f$. It follows that, for $\mathcal{H}^{n}$-a.e. $y\in\Omega^{\prime}$ we have that $f^{-1}(y)\cap B_{f}=\varnothing$. The co-area formula now yields
$\int_{\Omega^{\prime}}f_{\ast}g\mathrm{d}\mathcal{H}^{n}=\int_{\Omega^{\prime}}\sum_{x\in f^{-1}(y)}g(x)\mathrm{d}\mathcal{H}^{n}(y)=\int_{\Omega}g\mathbf{J}f\mathrm{d}\mathcal{H}^{n},$
as claimed. ∎
The main result of this subsection is the following curve-wise estimates for the multi-valued inverse. In the statement $K_{I}=K_{I}(f)$ and $K_{O}=K_{O}(f)$.
Theorem 3.7.
For $\operatorname{Mod}_{n}$-a.e. curve $\gamma$ in $f(\Omega)$, $\accentset{\leftarrow}{f}\circ\gamma$ is an absolutely continuous curve in $\mathcal{A}_{d}(\Omega)$, and satisfies
$H(\gamma_{t})|\gamma_{t}^{\prime}|\leq|(\accentset{\leftarrow}{f}\circ\gamma)^{\prime}_{t}|\leq(K_{I}K_{O})^{1/n}H(\gamma_{t})|\gamma_{t}^{\prime}|$ (3.3)
a.e. $t$, where
$H(y):=\big(f_{\ast}(\|Df\|^{-2})(y)\big)^{1/2},\quad y\in f(\Omega).$ (3.4)
Proof.
If $\Gamma_{not}$ is the family of paths in $\Omega$ which are not locally absolutely continuous, then $\operatorname{Mod}_{n}\Gamma_{not}=0$. By Poletski’s inequality [26, Chapter II, Theorem 8.1] we have that $\operatorname{Mod}_{n}f(\Gamma_{not})\leq K_{I}\operatorname{Mod}_{n}\Gamma_{not}=0$. Thus, for $\operatorname{Mod}_{n}$-a.e. $\gamma$ in $\Omega^{\prime}$, every lift $\tilde{\gamma}$ of $\gamma$ is locally absolutely continuous. By Proposition 3.3 we have $\accentset{\leftarrow}{f}\circ\gamma=\llbracket\tilde{\gamma}_{1},\ldots,\tilde{\gamma}_{d}\rrbracket$, where $\tilde{\gamma}_{j}$ are the lifts of $\gamma$ by $f$ (counted with multiplicity). It follows that
$d_{\mathcal{A}}(\accentset{\leftarrow}{f}(\gamma_{t}),\accentset{\leftarrow}{f}(\gamma_{s}))^{2}\leq\sum_{l}^{d}|\tilde{\gamma}_{l}(t)-\tilde{\gamma}_{l}(s)|^{2},\quad s,t\in[0,1]$
which implies that $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous whenever each lift $\tilde{\gamma}_{l}$ is absolutely continuous. In particular, $\accentset{\leftarrow}{f}\circ\gamma$ is absolutely continuous for $\operatorname{Mod}_{n}$-a.e. $\gamma$.
We now prove (3.3). Let $\Gamma_{0}$ be a path family in $\Omega$ with $\operatorname{Mod}_{n}\Gamma_{0}=0=\operatorname{Mod}_{n}f(\Gamma_{0})$ so that $\|Df\|$ is an upper gradient of $f$ along every $\gamma\notin\Gamma_{0}$. Observe that $f(B_{f})$ is closed since $f$ has finite degree and that $|f(B_{f})|=0$ by [26, Chapter II, Theorem 7.4]. If $B=B(y,r)\subset\Omega^{\prime}\setminus f(B_{f})$ is a ball such that $f^{-1}B=U_{1}\cup\cdots\cup U_{d}$, $U_{l}$ and $f:U_{l}\to B$ is quasiconformal, denote by $g_{l}:B\to U_{l}$ the inverse of $f|_{U_{l}}$, for $l=1,\ldots,d$. Then
$1\leq\|Df(g_{l}(y))\|^{n}\|Dg_{l}(y)\|^{n}\leq K_{O}\mathbf{J}f(g_{l}(y))K_{I}\mathbf{J}g(y)=K_{O}K_{I}$ (3.5)
a.e. $y\in B$, see [29]. By Vitali’s covering theorem there exist countably many disjoint balls $B_{j}=B(y_{j},r_{j})$ as above, with inverses $g^{j}_{l}:B_{j}\to U_{l}^{j}$ satisfying (3.5) outside a null set $N_{j}\subset B_{j}$, such that $\big|[\Omega^{\prime}\setminus f(B_{f})]\setminus\bigcup_{j}B_{j}\big|=0$. Denote $E=f(B_{f})\cup\Omega^{\prime}\setminus\bigcup_{j}B_{j}$, so that $|E|=0$. For each $j$, let $\Gamma^{j}$ be a path-family in $B_{j}$ with $\operatorname{Mod}_{n}\Gamma^{j}=0$ such that $\|Dg_{l}^{j}\|$ is an upper gradient along $\gamma$, and $|\gamma^{-1}(N_{j})|=0$, for every $\gamma\notin\Gamma^{j}$. Let $\Gamma_{1}$ by the family of paths in $\Omega^{\prime}$ which have a subcurve in some $\Gamma^{j}$. Then $\operatorname{Mod}_{n}\Gamma_{1}=0$. Finally, define $\Gamma_{2}=\Gamma_{1}\cup f(\Gamma_{0})\cup\Gamma_{E}^{+}$ and note that $\operatorname{Mod}_{n}\Gamma_{2}=0$.
Let $\gamma\notin\Gamma_{2}$. Then $[0,1]$ can be covered up to a null-set by open sets $\gamma^{-1}(B_{j})$, and for a.e. $t\in\gamma^{-1}(B_{j})$ we have
$|(g^{j}_{l}\circ\gamma)_{t}^{\prime}|$ $\leq\|Dg^{j}_{l}(\gamma_{t})\||\gamma_{t}^{\prime}|$
$|\gamma_{t}^{\prime}|=|(f\circ g_{l}^{j}\circ\gamma)_{t}^{\prime}|$ $\leq\|Df(g_{l}^{j}(\gamma_{t}))\||(g_{l}^{j}\circ\gamma)_{t}^{\prime}|$
$\frac{1}{\|Df(g_{l}^{j}(\gamma_{t}))\|}$ $\leq\|Dg^{j}_{l}(\gamma_{t})\|\leq\frac{(K_{I}K_{O})^{1/n}}{\|Df(g_{l}^{j}(\gamma_{t}))\|}.$
Thus
$\frac{|\gamma_{t}^{\prime}|}{\|Df(g_{l}^{j}(\gamma_{t}))\|}\leq|(g_{l}^{j}\circ\gamma)_{t}^{\prime}|\leq\frac{(K_{I}K_{O})^{1/n}|\gamma_{t}^{\prime}|}{\|Df(g_{l}^{j}(\gamma_{t}))\|}\quad\textrm{a.e. }t\in\gamma^{-1}(B_{j})$
Since $F\circ\gamma=\llbracket g_{1}^{j}\circ\gamma,\ldots,g_{d}^{j}\circ\gamma\rrbracket$ on $\gamma^{-1}(B_{j})$, we obtain
$|(\accentset{\leftarrow}{f}\circ\gamma)_{t}^{\prime}|^{2}$ $=\sum_{l}^{d}|(g_{l}^{j}\circ\gamma)^{\prime}_{t}|^{2}\leq(K_{I}K_{O})^{2/n}\sum_{l}^{d}\frac{|\gamma_{t}^{\prime}|^{2}}{\|Df(g_{l}^{j}(\gamma_{t}))\|^{2}}$
$=(K_{I}K_{O})^{2/n}f_{\ast}(\|Df\|^{-2})(\gamma_{t})|\gamma_{t}^{\prime}|^{2},$
and similarly
$|(\accentset{\leftarrow}{f}\circ\gamma)_{t}^{\prime}|^{2}\geq f_{\ast}(\|Df\|^{-2})(\gamma_{t})|\gamma_{t}^{\prime}|^{2}$
a.e. $t\in\gamma^{-1}(B_{j})$ for every $j$. This proves (3.3) and finishes the proof of the Proposition. ∎
Corollary 3.8.
The multivalued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\Omega)$ is Sobolev: $\accentset{\leftarrow}{f}\in W^{1,n}_{loc}(f(\Omega),\mathcal{A}_{d}(\Omega))$. More precisely, the minimal $n$-weak upper gradient $|D\accentset{\leftarrow}{f}|$ of $\accentset{\leftarrow}{f}$ satisfies
$H\leq|D\accentset{\leftarrow}{f}|\leq(K_{I}K_{O})^{1/n}H$
and
$\int_{U}|D\accentset{\leftarrow}{f}|^{n}\mathrm{d}y\leq d^{n/2-1}K_{I}K_{O}|f^{-1}U|$
for any compact $U\subset f(\Omega)$.
Proof.
By Theorem 3.7 (and using its notation), $(K_{I}K_{O})^{1/n}H$ is a weak upper gradient of $F$. Thus $|D\accentset{\leftarrow}{f}|\leq(K_{I}K_{O})^{1/n}H$. Using e.g. [3, Theorem 1.1] we obtain $H\leq|D\accentset{\leftarrow}{f}|$ from the left-hand inequality in (3.3). For any compact $U\subset f(\Omega)$ we have, by Lemma 3.6, that
$\int_{U}H^{n}\mathrm{d}y$ $\leq d^{n/2-1}\int_{U}f_{\ast}(\|Df\|^{-n})\mathrm{d}y$
$=d^{n/2-1}\int_{f^{-1}U}\frac{\mathbf{J}f}{\|Df\|^{n}}\mathrm{d}x\leq d^{n/2-1}|f^{-1}U|.$
This completes the proof of the claim. ∎
3.3. Area formula
We continue assuming $f:\Omega\to\mathbb{R}^{n}$ is a non-constant quasiregular map of degree $d$. We will prove that the multi-valued inverse satisfies Lusin’s properties $(N)$ and $(N^{-1})$ as well as the area formula.
Proposition 3.9.
The multi-valued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\mathcal{A}_{d}(\Omega)$ satisfies Lusin’s properties $(N)$ and $(N^{-1})$: Given $E\subset f(\Omega)$, we have
$\mathcal{H}^{n}(\accentset{\leftarrow}{f}(E))=0\Longleftrightarrow|E|=0.$
Moreover, $\accentset{\leftarrow}{f}$ is differentiable almost everywhere on $f(\Omega)$ and satisfies the area formula
$\int_{f(\Omega)}g\circ\accentset{\leftarrow}{f}\ \mathbf{J}\accentset{\leftarrow}{f}\mathrm{d}y=\int_{\Omega_{f}}g\mathrm{d}\mathcal{H}^{n},$ (3.6)
where $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega))\subset\mathcal{A}_{d}(\Omega)$.
Property $(N)$ and the area formula follow from the pseudomonotonicity of $\accentset{\leftarrow}{f}$ (see Lemma 3.10 below) by results in [10, 17], while the inverse Lusin property $(N^{-1})$ is a consequence of Proposition 3.4. Recall that a map $F:X\to Y$ is pseudomonotone if there exists a constants $K,r_{0}>0$ such that
$\operatorname{diam}F(B(x,r))\leq K\operatorname{diam}\partial F(B(x,r))$
for $x\in X$ and $r\in(0,r_{0})$.
Lemma 3.10.
Let $f:\Omega\to\mathbb{R}^{n}$ be a branched cover of degree $d$. Then for every compact $K\subset f(\Omega)$ there exists $r_{0}>0$ such that
$\operatorname{diam}F(B(y,r))\leq\sqrt{d}\operatorname{diam}\partial F(B(y,r))$
for all $y\in K$ and $r\in(0,r_{0})$.
Proof.
Let $r_{0}$ be the Lebesgue number of the open cover $\{B(y,r_{y})\}_{y\in K}$ where, for each $y\in K$, $r_{y}>0$ is such that
$f^{-1}B(y,r_{y})=\bigcup_{x\in f^{-1}(y)}U_{f}(x,r_{y})$
and each $U_{f}(x,r_{y})$ is a normal domain for $f$ compactly contained in $\Omega$. Now for each $y\in K$ and $r<r_{y}$ we have $B(y,r)\subset B(y_{0},r_{y_{0}})$ for some $y_{0}\in K$. Writing
$f^{-1}B(y,r)=\bigcup_{x\in f^{-1}(y)}U_{f}(x,r),$
we have
$\operatorname{diam}F(B(y,r))^{2}$ $\leq\sum_{x\in f^{-1}(y)}\iota(f,x)\operatorname{diam}U_{f}(x,r)^{2}\leq d\max_{x\in f^{-1}(y)}\operatorname{diam}U_{f}(x,r)^{2}$
$\leq d\max_{x\in f^{-1}(y)}\operatorname{diam}\partial U_{f}(x,r)^{2}\leq d\operatorname{diam}\partial F(B(y,r))^{2}.$
See [26, Chapter I, Lemma 4.7] for the fact that $U_{f}(x,r)$ is a normal domain of $f$ and thus $f(\partial U_{f}(x,r))=\partial B(y,r)$, implying
$\operatorname{diam}\partial U_{f}(x,r)\leq\operatorname{diam}F(\partial B(y,r))=\operatorname{diam}\partial F(B(y,r)).$
Proof of Proposition 3.9.
Since $\accentset{\leftarrow}{f}\in W^{1,n}_{loc}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ is continuous and pseudomonotone, it has Lusin’s property $(N)$ by [10, Theorem 7.2]. Moreover, by Corollary 3.4 in [17] $\accentset{\leftarrow}{f}$ is metrically differentiable almost everywhere, while by Corollary 3.9(1) in [17], $\accentset{\leftarrow}{f}$ satisfies the area formula (3.6).
The a.e. metric differentiability of $f$ implies that $\operatorname{Lip}f(x)<\infty$ a.e. $x\in f(\Omega)$, thus by the Stepanov theorem for multivalued maps (see [20, Corollary 2.8]) it follows that $f$ is differentiable almost everywhere.
It remains to prove Lusin’s inverse property $(N^{-1})$. This follows from Proposition 3.4 with $s=n$: if $E\subset f(\Omega)$ and $\mathcal{H}^{n}(\accentset{\leftarrow}{f}(E))=0$, then
$|f^{-1}(E)|=|(\accentset{\leftarrow}{f}\circ f)^{-1}(\accentset{\leftarrow}{f}(E))|\leq d\cdot\mathcal{H}^{n}(\accentset{\leftarrow}{f}(E))=0,$
which implies $|E|=0$ by the Lusin property $(N)$ of $f$. ∎
4. Pull-back of differential forms by multi-valued Sobolev maps
4.1. Differential forms on Almgren space
Let $\Gamma$ be a finite group acting by isometries on $\mathbb{R}^{N}$ and let $W\subset\mathbb{R}^{N}$ be a $\Gamma$-invariant open set. A smooth differential form $\omega\in\Omega^{k}(W)$ is called $\Gamma$-invariant, if $\gamma^{*}\omega=\omega$ for all $\gamma\in\Gamma$. $\Gamma$-invariant differential forms can be considered differential forms on $W/\Gamma$, even when this space is not a manifold. The particular case $\Gamma=S_{d}$ gives rise to differential forms on Almgren space.
Definition 4.1.
Suppose $V\subset\mathbb{R}^{n}$ is an open set, $d\geq 1$, and $k\in\mathbb{N}$. A (smooth) $k$-form on $\mathcal{A}_{d}(V)$ is a (smooth) $S_{d}$-invariant $k$-form $\omega\in\Omega^{k}(V^{d})$. We denote the space of smooth $k$-forms on $\mathcal{A}_{d}(V)$ by $\Omega^{k}(\mathcal{A}_{d}(V))$.
An element $\omega\in\Omega^{k}(\mathcal{A}_{d}(V))$ is a smooth map $\omega:V^{d}\to\bigwedge^{k}(\mathbb{R}^{n})^{d}$ with
$\omega_{x}(v_{1},\ldots,v_{k})=\omega_{\sigma(x)}(\sigma v_{1},\ldots,\sigma v_{k}),\quad x\in V^{d},\ v_{1},\ldots,v_{k}\in(\mathbb{R}^{n})^{d}.$ (4.1)
In particular, for each $\llbracket x\rrbracket\in\mathcal{A}_{d}(V)$, the value $\omega_{\llbracket x\rrbracket}:=\omega_{\sigma(x)}\circ\sigma\in\bigwedge^{k}(\mathbb{R}^{n})^{d}$ is well-defined and independent of $\sigma\in S_{d}$. We define the comass $\|\omega\|:\mathcal{A}_{d}(V)\to\mathbb{R}$ of a $k$-form $\omega\in\Omega^{k}(\mathcal{A}_{d}(V))$ by
$\|\omega\|_{\llbracket x\rrbracket}=\|\omega_{x}\|,\quad\llbracket x\rrbracket\in\mathcal{A}_{d}(V).$
This is well-defined by (4.1).
The next lemma provides a projection map from the space of differential forms to the space of invariant differential forms. Later on we will need Lemma 4.2 for $\Gamma=S_{d_{0}}\times S_{d_{1}}\leq S_{d_{0}+d_{1}}$ as well as $\Gamma=S_{d}$.
Lemma 4.2.
Let $\Gamma$ be a finite group acting by linear isometries on $\mathbb{R}^{N}$, and $W\subset\mathbb{R}^{N}$ a $\Gamma$-invariant open set. For any $\omega\in\Omega^{k}(W)$, define
$P_{\Gamma}\omega=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\gamma^{*}\omega.$
Then $P_{\Gamma}:\Omega^{k}(W)\to\Omega^{k}(W)$ is linear and continuous with respect to pointwise converge of forms. Moreover, for any $\omega\in\Omega^{k}(W)$,
  • (i)
    $P_{\Gamma}\omega$ is $\Gamma$-invariant,
  • (ii)
    $P_{\Gamma}\omega=\omega$ if and only if $\omega$ is $\Gamma$-invariant,
  • (iii)
    $P_{\Gamma}(\mathrm{d}\omega)=\mathrm{d}(P_{\Gamma}\omega)$,
  • (iv)
    $\|P_{\Gamma}\omega\|_{\infty}\leq\|\omega\|_{\infty}$.
Proof.
Linearity and continuity follow from the corresponding properties of the pull-back $\gamma^{*}\omega$. The identity
$\gamma_{0}^{*}(P_{\Gamma}\omega)=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\gamma_{0}^{*}\gamma^{*}\omega=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}(\gamma\circ\gamma_{0})^{*}\omega=P_{\Gamma}\omega,\quad\gamma_{0}\in\Gamma$
proves (i). To see (ii), note that if $\omega$ is $\Gamma$-invariant, then clearly $P_{\Gamma}\omega=\omega$, while if $P_{\Gamma}\omega=\omega$, then $\omega$ is $\Gamma$-invariant by (i).
Finally, (iii) follows from the fact that pull-back and exterior derivative commute:
$\mathrm{d}(P_{\Gamma}\omega)=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\mathrm{d}(\gamma^{*}\omega)=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\gamma^{*}(\mathrm{d}\omega)=P_{\Gamma}(\mathrm{d}\omega),$
and (iv) follows from the estimate
$\|P_{\Gamma}\omega_{x}\|\leq\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\|\gamma^{*}\omega\|_{\infty}\leq\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\|\omega\|_{\infty}\leq\|\omega\|_{\infty}.$
When $\Gamma=S_{d}$ acts on $(\mathbb{R}^{n})^{d}$ (see Section 2.2), we denote $\bar{\omega}:=P_{S_{d}}\omega$, and single out the following corollary.
Corollary 4.3.
The map $P=\omega\mapsto\bar{\omega}:\Omega^{k}(V^{d})\to\Omega^{k}(\mathcal{A}_{d}(V))$ is a linear projection onto $\Omega^{k}(\mathcal{A}_{d}(V))$ satisfying $\mathrm{d}(P\omega)=P(\mathrm{d}\omega)$. In particular, smooth $k$-forms on $\mathcal{A}_{d}(V)$ form a vector space and are closed under the exterior derivative.
Canonical $n$-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$
Let $\alpha$ be a smooth $k$-form on an open set $V\subset\mathbb{R}^{n}$ and consider the trace $\operatorname{tr}_{d}(\omega)=\operatorname{tr}(\omega)$, which is a $k$-form on the product $V^{d}$, defined by
$\operatorname{tr}(\alpha)=\sum_{j}^{d}P_{j}^{*}\alpha,$ (4.2)
where $P_{j}:(\mathbb{R}^{n})^{d}\to\mathbb{R}^{n}$ is the projection to the $j^{\textrm{th}}$ factor. It is straightforward to see from (4.2) that $\operatorname{tr}(\alpha)$ is a smooth $k$-form on $\mathcal{A}_{d}(V)$.
Definition 4.4.
We define a natural $n$-form $\omega_{n}$ on $\mathcal{A}_{d}(\mathbb{R}^{n})$ by $\omega_{n}:=\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}})$, where $\operatorname{vol}_{\mathbb{R}^{n}}$ is the volume form on $\mathbb{R}^{n}$.
Lemma 4.5.
We have that $\|\omega_{n}\|=1$ pointwise everywhere.
Proof.
Let $v_{l}=(v_{l}^{1},\ldots,v_{l}^{d})\in(\mathbb{R}^{n})^{d}$ for $l=1,\ldots,n$ be vectors with $|v_{l}|^{2}=\displaystyle\sum_{j}^{d}|v_{l}^{j}|^{2}\leq 1$. Then, using the Hadamard inequality and the generalized Hölder inequality, we may estimate
$\omega_{n}(v_{1},\ldots,v_{n})$ $=\sum_{j}^{d}v_{1}^{j}\wedge\cdots\wedge v_{n}^{j}\leq\sum_{j}^{d}|v_{1}^{j}|\cdots|v_{n}^{j}|$
$\leq\Big(\sum_{j}^{d}|v_{1}^{j}|^{n}\Big)^{1/n}\cdots\Big(\sum_{j}^{d}|v_{n}^{j}|^{n}\Big)^{1/n}\leq 1.$
Thus $\|\omega_{n}\|\leq 1$ everywhere. On the other hand, choosing $v_{l}=(e_{l},0,\ldots,0)$ yields $\omega_{n}(v_{1},\ldots,v_{n})=e_{1}\wedge\cdots\wedge e_{n}=1$, implying $\|\omega_{n}\|\geq 1$. This completes the proof. ∎
4.2. Pull-back by multi-valued maps
The purpose of this section is to define pull-backs of forms on $\mathcal{A}_{d}(\mathbb{R}^{n})$ by multi-valued maps, and to prove Theorem 1.1. The Sobolev regularity of pull-backs is interesting in its own right, and is also necessary for the proof of Corollary 1.3 and other applications. We start by defining the pull-back.
Suppose $U\subset\mathbb{R}^{m}$ and $V\subset\mathbb{R}^{n}$ are open, $h:U\to\mathcal{A}_{d}(V)$ is approximately differentiable at $x\in U$, and $h(x)=\llbracket y_{1},\ldots,y_{d}\rrbracket$ and $D_{x}h=\llbracket D_{x}h_{1},\ldots,D_{x}h_{d}\rrbracket$ are labelings such that (i) and (ii) in Theorem 2.2 hold. By (4.1), the element
$\omega_{y_{1},\ldots,y_{d}}\circ(D_{x}h_{1},\ldots,D_{x}h_{d})\in\bigwedge^{k}\mathbb{R}^{m}$
is independent of the permutation, i.e.
$\omega_{y_{\sigma^{-1}(1)},\ldots,y_{\sigma^{-1}(d)}}\circ(D_{x}h_{\sigma^{-1}(1)},\ldots,D_{x}h_{\sigma^{-1}(d)})=\omega_{y_{1},\ldots,y_{d}}\circ(D_{x}h_{1},\ldots,D_{x}h_{d})$
for all $\sigma\in S_{d}.$ We denote this common element of $\bigwedge^{k}\mathbb{R}^{m}$ by $\omega_{h(x)}\circ D_{x}h$.
Definition 4.6.
Suppose $U\subset\mathbb{R}^{m}$ and $V\subset\mathbb{R}^{n}$ are open sets. Let $\omega$ be a smooth $k$-form on $\mathcal{A}_{d}(V)$ and $h:U\to\mathcal{A}_{d}(V)$ approximately differentiable a.e. in $U$. We define a $k$-form $h^{*}\omega:U\to\bigwedge^{k}\mathbb{R}^{m}$ (defined almost everywhere) by
$(h^{*}\omega)_{x}=\omega_{h(x)}\circ D_{x}h.$
whenever $h$ is approximately differentiable at $x\in U$.
We now state the main result in the proof of Theorem 1.1, which states that pull-backs by multi-valued Lipschitz maps are flat forms.
Theorem 4.7.
Suppose $U\subset\mathbb{R}^{m}$ is an open set, $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is Lipschitz and $\omega\in\Omega^{k}(\mathcal{A}_{d}(\mathbb{R}^{n}))$. Then $\mathrm{d}(f^{*}\omega)=f^{*}(\mathrm{d}\omega)$ weakly. In particular, $f^{*}\omega$ is a flat form on $U$ in the sense of Whitney.
Theorem 1.1 follows from Theorem 4.7 by a Lusin-type approximation, cf. Proposition 2.3. The proof of Theorem 4.7 is more involved; it uses local decompositions of multi-valued functions, the density of tensor products of differential forms, and careful approximation arguments.
We begin with a simple estimate of the comass of the pull-back.
Lemma 4.8.
Suppose $U\subset\mathbb{R}^{m}$ and $V\subset\mathbb{R}^{n}$ are open sets, $\omega$ is a smooth $k$-form on $\mathcal{A}_{d}(V)$ and $h:U\to\mathcal{A}_{d}(V)$ a.e. approximately differentiable. Then $h^{*}\omega$ is a measurable $k$-form on $U$ and
$\|h^{*}\omega\|_{x}\leq\|D_{x}h\|_{x}^{k}\|\omega_{h(x)}\|\quad\textrm{a.e. }x\in U.$
Remark 4.9.
In particular $h^{*}\omega\in L^{p/k}_{loc}(U)$ if $h\in W^{1,p}_{loc}(U,\mathcal{A}_{d}(V))$.
Proof.
The approximate differentials $Dh_{1},\ldots,Dh_{d}$ are measurable. For all $x\in U$ where $h$ is approximately differentiable, we have
$\|h^{*}\omega\|_{x}$ $=\max\{\omega_{h(x)}(Dh(v_{1}),\ldots,Dh(v_{d})):\ |v_{1}|,\ldots,|v_{d}|\leq 1\}$
$\leq\|D_{x}h\|_{op}^{k}\|\omega_{h(x)}\|,$
completing the proof. ∎
For the next lemma, recall the definition of the tensor product of differential forms. For (open) manifolds $U_{0},U_{1}$, the projections $P_{i}:U_{0}\times U_{1}\to U_{i}$ induce linear maps $P_{i}^{*}:\Omega^{l}(U_{i})\to\Omega^{l}(U_{0}\times U_{1})$ which can be combined to obtain a bi-linear map
$P^{*}:\Omega^{l}(U_{0})\times\Omega^{r}(U_{1})\to\Omega^{k}(U_{0}\times U_{1}),\quad(\omega_{0},\omega_{1})\mapsto P_{0}^{*}\omega_{0}\wedge P_{1}^{*}\omega_{1}$
whenever $l+r=k$. The associated linear map $P^{*}:\Omega^{l}(U_{0})\otimes\Omega^{r}(U_{1})\to\Omega^{k}(U_{0}\times U_{1})$ is injective, and
$P^{*}:\bigcup_{r+l=k}\Omega^{l}(U_{0})\otimes\Omega^{r}(U_{1})\to\Omega^{k}(U_{0}\times U_{1})$
has dense range in the natural Frechet topology on $\Omega^{k}(U_{0}\times U_{1})$. We denote $\omega_{0}\otimes\omega_{1}:=P^{*}(\omega_{0},\omega_{1})$.
Note that, if $\omega_{i}\in\Omega^{k_{i}}(\mathcal{A}_{d_{i}}(\mathbb{R}^{n}))$ for $i=0,1$, and we denote $k=k_{0}+k_{1}$, $d=d_{0}+d_{1}$, the tensor product $\omega_{0}\otimes\omega_{1}\in\Omega^{k}((\mathbb{R}^{n})^{d})$ is not $S_{d}$-invariant but merely $S_{d_{0}}\times S_{d_{1}}$-invariant.
Lemma 4.10.
Let $\omega_{i}\in\Omega^{k_{i}}(V^{d_{i}})$, $i=0,1$, and $k=k_{0}+k_{1}$, $d=d_{0}+d_{1}$. Then
$P_{S_{d_{0}}\times S_{d_{1}}}(\omega_{0}\otimes\omega_{1})=(P_{S_{d_{0}}}\omega_{0})\otimes(P_{S_{d_{1}}}\omega_{1}).$ (4.3)
In particular, if $\omega_{i}$ is a $k_{i}$-form on $\mathcal{A}_{d_{i}}(V)$ for each $i=0,1$, then $\omega_{0}\otimes\omega_{1}$ is $S_{d_{0}}\times S_{d_{1}}$-invariant.
Proof.
The last claim follows directly from (4.3) and Lemma 4.2. To show (4.3), let $(\sigma_{0},\sigma_{1})\in S_{d_{0}}\times S_{d_{1}}$. For any $v\in(\mathbb{R}^{n})^{d}$, $(\sigma_{0},\sigma_{1})v=:v^{(\sigma_{0},\sigma_{1})}$ satisfies $P_{0}(v^{(\sigma_{0},\sigma_{1})})=\sigma_{0}(P_{0}v)$ and $P_{1}(v^{(\sigma_{0},\sigma_{1})})=\sigma_{1}(P_{1}v)$. Thus
$(\sigma_{0},\sigma_{1})^{*}(P_{0}^{*}\omega_{0}$ $\wedge P_{1}^{*}\omega_{1})(v_{1},\ldots,v_{k})=(P_{0}^{*}\omega_{0}\wedge P_{1}^{*}\omega_{1})(v_{1}^{(\sigma_{0},\sigma_{1})},\ldots,v_{k}^{(\sigma_{0},\sigma_{1})})$
$=\frac{1}{k!}\sum_{\theta\in S_{k}}\operatorname{sgn}(\theta)$ $(P_{0}^{*}\omega_{0})(v_{\theta(1)}^{(\sigma_{0},\sigma_{1})},\ldots,v_{\theta(k_{0})}^{(\sigma_{0},\sigma_{1})})(P_{1}^{*}\omega_{1})(v_{\theta(k_{0}+1)}^{(\sigma_{0},\sigma_{1})},\ldots,v_{\theta(k)}^{(\sigma_{0},\sigma_{1})})$
$=\frac{1}{k!}\sum_{\theta\in S_{k}}\operatorname{sgn}(\theta)$ $\omega_{0}(\sigma_{0}(P_{0}v_{\theta(1)}),\ldots,\sigma_{0}(P_{0}v_{\theta(k_{0})}))$
$\times\omega_{1}(\sigma_{1}(P_{1}v_{\theta(k_{0}+1)}),\ldots,\sigma_{1}(P_{1}v_{\theta(k)}))$
$=P_{0}^{*}(\sigma_{0}^{*}\omega_{0})\wedge P_{1}$ $^{*}(\sigma_{1}^{*}\omega_{1})(v_{1},\ldots,v_{k})$
for all $v_{1},\ldots,v_{k}\in(\mathbb{R}^{n})^{d}$. Now we can calculate
$P_{S_{d_{0}}\times S_{d_{1}}}(\omega_{0}\otimes\omega_{1})=$ $P_{S_{d_{0}}\times S_{d_{1}}}(P_{0}^{*}\omega_{0}\wedge P_{1}^{*}\omega_{1})$
$=$ $\frac{1}{|S_{d_{0}}||S_{d_{1}}|}\sum_{\sigma_{0}\in S_{d_{0}}}\sum_{\sigma_{1}\in S_{d_{1}}}(\sigma_{0},\sigma_{1})^{*}(P_{0}^{*}\omega_{0}\wedge P_{1}^{*}\omega_{1})$
$=$ $\frac{1}{|S_{d_{0}}||S_{d_{1}}|}\sum_{\sigma_{0}\in S_{d_{0}}}\sum_{\sigma_{1}\in S_{d_{1}}}P_{0}^{*}(\sigma_{0}^{*}\omega_{0})\wedge P_{1}^{*}(\sigma_{1}^{*}\omega_{1})$
$=$ $P_{0}^{*}(P_{S_{d_{0}}}\omega_{0})\wedge P_{1}^{*}(P_{S_{d_{1}}}\omega_{1})=(P_{S_{d_{0}}}\omega_{0})\otimes(P_{S_{d_{1}}}\omega_{1}).$
In the next lemma we show that $S_{d_{0}}\times S_{d_{1}}$-invariant forms have a well-defined pull back by multivalued functions $f=\llbracket f_{0},f_{1}\rrbracket$, where $f_{i}:U\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$ is Lipschitz for each $i=0,1$. We emphasize that $S_{d_{0}}\times S_{d_{1}}$-invariant forms cannot be pulled back by general multivalued Lipchitz maps $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$, that is, we are using the additional information of having a decomposition in Lemma 4.11.
Lemma 4.11.
Suppose $f_{i}:U\to\mathcal{A}_{d_{i}}(V)$ is Lipschitz ($i=0,1$), $d=d_{0}+d_{1}$, and $\omega$ is a $S_{d_{0}}\times S_{d_{1}}$-invariant form on $V^{d}$. Then the pull-back $\llbracket f_{0},f_{1}\rrbracket^{*}\omega$ is a well-defined $k$-form on $U$. Moreover,
$\llbracket f_{0},f_{1}\rrbracket^{*}(\omega_{0}\otimes\omega_{1})=f_{0}^{*}\omega_{0}\wedge f_{1}^{*}\omega_{1}$
whenever $\omega_{i}\in\Omega^{k_{i}}(\mathcal{A}_{d_{i}}(V))$ ($i=0,1$) and $k=k_{0}+k_{1}$.
Proof.
Let $x\in U$ be such that $f_{0},f_{1}$ are differentiable at $x$, and let $f_{i}^{1}(x),\ldots,f_{i}^{d_{i}}(x)$ and $D_{x}f_{i}^{1},\ldots,D_{x}f_{i}^{d_{i}}$ be as in Theorem 2.2 (i) and (ii). (In particular, note that $f=\llbracket f_{0},f_{1}\rrbracket$ is differentiable at $x$ and $Df=\llbracket D_{x}f_{0}^{1},\ldots,D_{x}f_{0}^{d_{0}},D_{x}f_{1}^{1},\ldots,D_{x}f_{1}^{d_{1}}\rrbracket$.) Given $(\sigma_{0},\sigma_{1})\in S_{d_{0}}\times S_{d_{1}}$, denote
$D_{x}f^{(\sigma_{0},\sigma_{1})}=(D_{x}f_{0}^{\sigma_{0}^{-1}(1)},\ldots D_{x}f_{0}^{\sigma_{0}^{-1}(d_{0})},D_{x}f_{1}^{\sigma_{1}^{-1}(1)},\ldots,D_{x}f_{1}^{\sigma_{1}^{-1}(d_{1})})$
and similarly
$(f_{0}^{\sigma_{0}}(x),f_{1}^{\sigma_{1}}(x))=(f_{0}^{\sigma_{0}^{-1}(1)}(x),\ldots,f_{0}^{\sigma_{0}^{-1}(d_{0})},f_{1}^{\sigma_{1}^{-1}(1)},\ldots,f_{1}^{\sigma_{1}^{-1}(d_{1})}).$
We define
$(f^{*}\omega)_{x}(v_{1},\ldots,v_{k})=\omega_{(f_{0}^{\sigma_{0}}(x),f_{1}^{\sigma_{1}}(x))}(D_{x}f^{(\sigma_{0},\sigma_{1})}v_{1},\ldots,D_{x}f^{(\sigma_{0},\sigma_{1})}v_{k})$
for $v_{1},\ldots,v_{k}\in T_{x}U=\mathbb{R}^{m}$. By the $S_{d_{0}}\times S_{d_{1}}$-invariance of $\omega$, this expression is independent of $(\sigma_{0},\sigma_{1})$ and gives a well-defined element in $\bigwedge^{k}\mathbb{R}^{m}$, which we denote by $(f^{*}\omega)_{x}=\omega_{(f_{0}(x),f_{1}(x))}\circ(D_{x}f_{0},D_{x}f_{1})$.
In particular, if $\omega_{i}\in\Omega^{k_{i}}(\mathcal{A}_{d_{i}}(V))$, $i=0,1$, then $\omega_{0}\otimes\omega_{1}$ is $S_{d_{0}}\times S_{d_{1}}$-invariant. We may calculate
$f^{*}(\omega_{0}\otimes\omega_{1})(v_{1},\ldots,v_{k})=(\omega_{0}\otimes\omega_{1})((D_{x}f_{0},D_{x}f_{1})v_{1},\ldots,(D_{x}f_{0},D_{x}f_{1})v_{k})$
$=$ $\frac{1}{k!}\sum_{\theta\in S_{k}}\operatorname{sgn}(\theta)\omega_{0}(Df_{0}v_{\theta(1)},\ldots,Df_{0}v_{\theta(k_{0})})\omega_{1}(Df_{1}v_{\theta(k_{0}+1)},\ldots,Df_{1}v_{\theta(k)})$
$=$ $(f_{0}^{*}\omega_{0})\wedge(f_{1}^{*}\omega_{1})(v_{1},\ldots,v_{k})$
for every $v_{1},\ldots,v_{k}\in\mathbb{R}^{m}$. This completes the proof. ∎
Proposition 4.12.
Let $d_{0},d_{1}\geq 1$ be natural numbers, $d:=d_{0}+d_{1}$, and let $U\subset\mathbb{R}^{m}$ be an open set. If $f_{i}:U\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$ are Lipschitz functions such that $\mathrm{d}(f_{i}^{*}\omega)=f_{i}^{*}(\mathrm{d}\omega)$ weakly for all $l$-forms $\omega\in\Omega^{l}(\mathcal{A}_{d_{i}}(\mathbb{R}^{n}))$, $l\leq k$ ($i=0,1$), then $\mathrm{d}(\llbracket f_{0},f_{1}\rrbracket^{*}\omega)=\llbracket f_{0},f_{1}\rrbracket^{*}(\mathrm{d}\omega)$ weakly for all $k$-forms $\omega\in\Omega^{k}(\mathcal{A}_{d}(\mathbb{R}^{n}))$.
Proof.
Denote $f:=\llbracket f_{0},f_{1}\rrbracket$, and let $\omega\in\Omega^{k}(\mathcal{A}_{d}(\mathbb{R}^{n}))$. Let
$(\omega_{j})\subset\bigcup_{r+l=k}\Omega^{l}((\mathbb{R}^{n})^{d_{0}})\otimes\Omega^{r}((\mathbb{R}^{n})^{d_{1}})$
be a sequence such that $(\omega_{j})_{x}\to\omega_{x}$ and $(\mathrm{d}\omega_{j})_{x}\to(\mathrm{d}\omega)_{x}$ for all $x\in(\mathbb{R}^{n})^{d}$ and $\sup_{j}\max\{\|\omega_{j}\|_{\infty},\|\mathrm{d}\omega_{j}\|_{\infty}\}<\infty$. Then
$\tilde{\omega}_{j}$ $:=P_{S_{d_{0}}\times S_{d_{1}}}\omega_{j}\to P_{S_{d_{0}}\times S_{d_{1}}}\omega=\omega,$
$\mathrm{d}\tilde{\omega}_{j}$ $=P_{S_{d_{0}}\times S_{d_{1}}}(\mathrm{d}\omega_{j})\to P_{S_{d_{0}}\times S_{d_{1}}}(\mathrm{d}\omega)=\mathrm{d}\omega$
and $\sup_{j}\max\{\|\tilde{\omega}_{j}\|_{\infty},\|\mathrm{d}\tilde{\omega}_{j}\|_{\infty}\}<\infty$, cf. Lemma 4.2. By Lemma 4.11 each $\tilde{\omega}_{j}$ is a sum of tensor products $\omega_{0}\otimes\omega_{1}$ where $\omega_{i}$ is $S_{d_{i}}$-invariant and $f^{*}(\omega_{0}\otimes\omega_{1})=(f_{0}^{*}\omega_{0})\wedge(f^{*}_{1}\omega_{1})$. Together with the assumption that $f_{i}^{*}(\mathrm{d}\omega_{i})=\mathrm{d}(f^{*}_{i}\omega)$ weakly for $S_{d_{i}}$-invariant forms $\omega$, this yields
$f^{*}(\mathrm{d}(\omega_{0}\otimes\omega_{1}))$ $=f^{*}(\mathrm{d}\omega_{0}\otimes\omega_{1}+(-1)^{\deg\omega_{0}}\omega_{0}\otimes\mathrm{d}\omega_{1})$
$=f^{*}_{0}(\mathrm{d}\omega_{0})\wedge(f^{*}_{1}\omega_{1})+(-1)^{\deg\omega_{0}}(f_{0}^{*}\omega_{0})\wedge(f_{1}^{*}(\mathrm{d}\omega_{1}))$
$=\mathrm{d}(f_{0}^{*}\omega_{0})\wedge(f_{1}^{*}\omega_{1})+(-1)^{\deg\omega_{0}}(f_{0}^{*}\omega_{0})\wedge\mathrm{d}(f_{1}^{*}\omega_{1})$
$=\mathrm{d}(f_{0}^{*}\omega_{0}\wedge f_{1}^{*}\omega_{1})=\mathrm{d}f^{*}(\omega_{0}\otimes\omega_{1})$
weakly. Consequently $f^{*}(\mathrm{d}\tilde{\omega}_{j})=\mathrm{d}(f^{*}\tilde{\omega}_{j})$ weakly. Now the pointwise convergence $f^{*}\tilde{\omega}_{j}\to f^{*}\omega$, $f^{*}(\mathrm{d}\tilde{\omega}_{j})\to f^{*}(\mathrm{d}\omega)$ together with the dominated convergence theorem yield
$\int_{U}\alpha\wedge f^{*}(\mathrm{d}\omega)$ $=\lim_{j\to\infty}\int_{U}\alpha\wedge f^{*}(\mathrm{d}\tilde{\omega}_{j})=(-1)^{k+1}\lim_{j\to\infty}\int_{U}\mathrm{d}\alpha\wedge f^{*}\tilde{\omega}_{j}$
$=(-1)^{k+1}\int_{U}\mathrm{d}\alpha\wedge f^{*}\omega$
for all $\alpha\in\Omega_{c}^{m-k-1}(U)$. This proves the claim. ∎
The next proposition gives uniform control of the Lipschitz constants of the approximations.
Proposition 4.13.
Let $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ be $L$-Lipschitz. Given $\varepsilon>0$, denote
$F_{\varepsilon}=\{x\in U:d_{\mathcal{A}}(f(x),d\llbracket b(f(x))\rrbracket<\varepsilon\},\quad\eta_{\varepsilon}=\Big(1-\frac{\operatorname{dist}(F_{\varepsilon},\cdot)}{\varepsilon}\Big)_{+}$
Define $f_{\varepsilon}:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ by
$f_{\varepsilon}=\llbracket(1-\eta_{\varepsilon})f_{1}+\eta_{\varepsilon}b(f),\ldots,(1-\eta_{\varepsilon})f_{d}+\eta_{\varepsilon}b(f)\rrbracket,$
where $f=\llbracket f_{1},\ldots,f_{d}\rrbracket$. Then $f_{\varepsilon}\to f$ locally uniformly and $\operatorname{LIP}(f_{\varepsilon})\leq(3+2d)L$.
Proof.
Denote $g_{j}=b(f)-f_{j}$ and note that $f_{\varepsilon}=\llbracket f_{1}+\eta_{\varepsilon}g_{1},\ldots,f_{d}+\eta_{\varepsilon}g_{d}\rrbracket$. Observe moreover that (a) $\eta_{\varepsilon}$ is $\varepsilon^{-1}$-Lipschitz, (b) $\eta_{\varepsilon}|_{F_{\varepsilon}}=1$, and (c) $\operatorname{spt}(\eta_{\varepsilon})\subset B(F_{\varepsilon},\varepsilon)\subset\overline{F}_{(1+L)\varepsilon}$. We may assume $L\geq 1$ so that (c) implies $\operatorname{spt}(\eta_{\varepsilon})\subset F_{2L\varepsilon}$.
Fix $x,y\in U$. For any $\sigma\in S_{d}$ we have
$\Big(\sum_{j}^{d}|f_{j}(x)-f_{\sigma(j)}(y)+\eta_{\varepsilon}(x)g_{j}(x)-\eta_{\varepsilon}(y)g_{\sigma(j)}(y)|^{2}\Big)^{1/2}$
$\leq$ $\Big(\sum_{j}^{d}|f_{j}(x)-f_{\sigma(j)}(y)|^{2}\Big)^{1/2}+\Big(\sum_{j}^{d}|\eta_{\varepsilon}(x)g_{j}(x)-\eta_{\varepsilon}(y)g_{\sigma(j)}(y)|^{2}\Big)^{1/2}$
If $x,y\notin F_{2L\varepsilon}$, then $|\eta_{\varepsilon}(x)g_{j}(x)-\eta_{\varepsilon}(y)g_{\sigma(j)}(y)|=0$, while if $x\in F_{2L\varepsilon}$ or $y\in F_{2L\varepsilon}$, then $\min\{|g_{j}(x)|,g_{\sigma(j)}(y)\}\leq 2L\varepsilon$ and $\max\{\eta_{\varepsilon}(x),\eta_{\varepsilon}(y)\}\leq 1$. Consequently
$|\eta_{\varepsilon}(x)g_{j}(x)-\eta_{\varepsilon}(y)g_{\sigma(j)}(y)|\leq$ $|\eta_{\varepsilon}(x)-\eta_{\varepsilon}(y)|2L\varepsilon+|g_{j}(x)-g_{\sigma(j)}(y)|$
$\leq$ $2Ld(x,y)+|f_{j}(x)-f_{\sigma(j)}(y)|+|b(f(x))-b(f(y))|$
$\leq$ $2L|x-y|+\frac{1}{\sqrt{d}}d_{\mathcal{A}}(f(x),f(y))+|f_{j}(x)-f_{\sigma(j)}(y)|.$
Thus we obtain the estimate
$\Big(\sum_{j}^{d}|(f_{\varepsilon})_{j}(x)-(f_{\varepsilon})_{\sigma(j)}(y)|^{2}\Big)^{1/2}$ $\leq 2\Big(\sum_{j}^{d}|f_{j}(x)-f_{\sigma(j)}(y)|^{2}\Big)^{1/2}$
$+$ $2Ld\cdot|x-y|+d_{\mathcal{A}}(f(x),f(y)).$
Taking infimum over $\sigma\in S_{d}$ yields
$d_{\mathcal{A}}(f_{\varepsilon}(x),f_{\varepsilon}(y))$ $\leq 3d_{\mathcal{A}}(f(x),f(y))+2Ld|x-y|$
$\leq(3+2d)L|x-y|.$
It remains to show that $f_{\varepsilon}\to f$ locally uniformly in $U$. Since $\operatorname{LIP}(f_{\varepsilon})$ is bounded uniformly in $\varepsilon$, it suffices to prove pointwise convergence. To this end observe that
$d_{\mathcal{A}}(f_{\varepsilon}(x),f(x))$ $\leq\eta_{\varepsilon}(x)\Big(\sum_{j}^{d}|f_{j}(x)-b(f(x))|^{2}\Big)^{1/2}$
$=\eta_{\varepsilon}(x)d_{\mathcal{A}}(f(x),d\llbracket b(f(x)\rrbracket)\leq 2L\varepsilon.$
Indeed, if $x\notin F_{2L\varepsilon}$ then $\eta_{\varepsilon}(x)=0$ and otherwise $d_{\mathcal{A}}(f(x),b(f(x))\leq 2L\varepsilon$. This completes the proof. ∎
Proof of Theorem 4.7.
We prove the claim by induction on $d$. For $d=1$ the result is classical. Suppose $d>1$ and that $g^{*}(\mathrm{d}\omega)=\mathrm{d}(g^{*}\omega)$ weakly for any locally Lipschitz map $g:U^{\prime}\to\mathcal{A}_{l}(\mathbb{R}^{n})$ from an open set $U^{\prime}\subset\mathbb{R}^{m}$ with $l<d$. We shall decompose the original map $f$ into sums of locally simpler maps and utilize Proposition 4.12.
Denote $F:=f^{-1}(\operatorname{diag}\mathcal{A}_{d}(\mathbb{R}^{n}))$ and $L=\operatorname{LIP}(f)$. Observe that $F$ is a closed set in $U$, and that $f=d\llbracket b(f)\rrbracket$ on $F$.
Given $\varepsilon>0$, set
$F_{\varepsilon}=\{x:d_{\mathcal{A}}(f(x),b(f(x))<\varepsilon\}\quad\textrm{and}\quad\eta_{\varepsilon}(y)=(1-\varepsilon^{-1}\operatorname{dist}(F_{\varepsilon},y))_{+}$
as in Proposition 4.13. Recall that $F_{\varepsilon}$ is open and $\bigcap_{\varepsilon>0}F_{\varepsilon}=F$, and that (a) $\eta_{\varepsilon}$ is $\varepsilon^{-1}$-Lipschitz, (b) $\eta_{\varepsilon}|_{F_{\varepsilon}}=1$, and (c) $\operatorname{spt}(\eta_{\varepsilon})\subset B(F_{\varepsilon},\varepsilon)\subset\overline{F}_{(1+L)\varepsilon}$. The map $f_{\varepsilon}:U\to\mathcal{A}_{d}(\mathbb{R}^{m})$
$f_{\varepsilon}=\llbracket(1-\eta_{\varepsilon})f_{1}+\eta_{\varepsilon}b(f),\ldots,(1-\eta_{\varepsilon})f_{d}+\eta_{\varepsilon}b(f)\rrbracket$
of Proposition 4.13 satisfies the following.
  • (0)
    $f_{\varepsilon}$ is $(3+2d)L$-Lipschitz;
  • (1)
    $f_{\varepsilon}|_{F}=f|_{F}$ and $f_{\varepsilon}|_{U\setminus F_{2L\varepsilon}}=f|_{U\setminus F_{2L\varepsilon}}$;
  • (2)
    $f_{\varepsilon}=d\llbracket b(f)\rrbracket$ on $F_{\varepsilon}$;
  • (3)
    For each $x\in U$ there exists a ball $B_{x}=B(x,r_{x})\subset U$ such that has a decomposition $f_{\varepsilon}|_{B_{x}}=\llbracket f_{\varepsilon,x}^{0},f_{\varepsilon,x}^{1}\rrbracket$.
Indeed, (0) follows from Proposition 4.13, and (1)-(2) are immediate consequences of the definition. We explain (3): given $x\in U$, if $x\in F_{\varepsilon}$, then the claim follows from (2). If $x\in U\setminus F_{\varepsilon}$ then in particular $x\in U\setminus F$ so that $|f_{i}(x)-f_{j}(x)|>0$ for some $i,j\in\{1,\ldots,d\}$. Now for these indices
$|(f_{\varepsilon})_{i}(x)-(f_{\varepsilon})_{i}(x)|=(1-\eta_{\varepsilon}(x))|f_{i}(x)-f_{j}(x)|>0.$
By [2, Proposition 1.6] there exists a ball $B_{x}=B(x,r_{x})\subset U$ and Lipschitz functions $f_{\varepsilon,x}^{i}:B_{x}\to\mathcal{A}_{d_{i}}(\mathbb{R}^{n})$ ($i=0,1$) with $d_{i}<d$, $d_{0}+d_{1}=d$, $\operatorname{LIP}(f_{x}^{i})\leq CL$ and $f_{\varepsilon}=\llbracket f_{\varepsilon,x}^{0},f_{\varepsilon,x}^{1}\rrbracket$ on $B_{x}$.
Fix a compact set $K\subset U$, and let $\{B_{x_{1}},\ldots,B_{x_{M}}\}$ be a finite cover of $\{B_{x}\}_{x\in K}$ where $B_{x}$ is as in (3) for each $x$. Let $\{\varphi_{1},\ldots\varphi_{M}\}$ be a partition of unity subordinate to $\{B_{x_{1}},\ldots,B_{x_{M}}\}$. We may express $f_{\varepsilon}^{*}\omega$ and $f_{\varepsilon}^{*}(\mathrm{d}\omega)$ as
$f_{\varepsilon}^{*}\omega=\sum_{l}^{M}\varphi_{l}\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}\omega,\quad f_{\varepsilon}^{*}(\mathrm{d}\omega)=\sum_{l}^{M}\varphi_{l}\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}(\mathrm{d}\omega)\quad\textrm{on }K.$
Thus for any $\alpha\in\Omega^{m-k-1}(U)$ supported in $K$ Proposition 4.12 yields
$\int\mathrm{d}\alpha\wedge f_{\varepsilon}^{*}\omega=$ $\sum_{l}^{M}\int\mathrm{d}\alpha\wedge(\varphi_{l}\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}\omega)=\sum_{l}^{M}\int\varphi_{l}\mathrm{d}\alpha\wedge\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}\omega$
$=$ $\sum_{l}^{M}\int\mathrm{d}(\varphi_{l}\alpha)\wedge\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}\omega$
$=$ $(-1)^{k+1}\sum_{l}^{M}\int\varphi_{l}\alpha\wedge\mathrm{d}(\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}\omega)$
$=$ $(-1)^{k+1}\sum_{l}^{M}\int\varphi_{l}\alpha\wedge\llbracket f_{\varepsilon,x_{l}}^{0},f_{\varepsilon,x_{l}}^{1}\rrbracket^{*}(\mathrm{d}\omega)=(-1)^{k+1}\int\alpha\wedge f_{\varepsilon}^{*}\omega.$
The equality in the second line above follows since $\sum_{l}^{M}\varphi_{l}=1$ on $K$. Since $K$ is arbitrary we obtain
$\int_{U}\mathrm{d}\alpha\wedge f_{\varepsilon}^{*}\omega=(-1)^{k+1}\int_{U}\alpha\wedge f_{\varepsilon}^{*}(\mathrm{d}\omega)$ (4.4)
for all $\alpha\in\Omega_{c}^{m-k-1}(U)$.
But by (2) we have that $f_{\varepsilon}^{*}\gamma=f^{*}\gamma$ almost everywhere outside $F_{2L\varepsilon}\setminus F$ for any differential form $\gamma$ on $\mathcal{A}_{d}(\mathbb{R}^{n})$. Together with (0) this yields
$\Big|\int_{U}\beta\wedge(f_{\varepsilon}^{*}\gamma-f^{*}\gamma)\Big|\leq C\|\beta\|_{\infty}L^{\deg(\gamma)}\|\gamma\|_{\infty}|F_{2L\varepsilon}\setminus F|\stackrel{{\scriptstyle\varepsilon\to 0}}{{\longrightarrow}}0.$
for any $\beta\in\Omega_{c}^{m-\deg(\gamma)}(U)$. Consequently
[…]
for all $\alpha\in\Omega_{c}^{m-k-1}(U)$. This completes the proof. ∎
Proof of Theorem 1.1.
Note that under the assumptions in the claim, we have $f^{*}\omega\in L^{p/k}_{loc}(U,\bigwedge^{k}\mathbb{R}^{m})$ and $f^{*}(\mathrm{d}\omega)\in L^{\frac{p}{k+1}}_{loc}(U,\bigwedge^{k+1}\mathbb{R}^{m})$ by Lemma 4.8. By considering an exhaustion of $U$ by relatively compact domains, we may assume that $U$ is bounded and that $f\in W^{1,p}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$. In particular $f\in W^{1,k}(U,\mathcal{A}_{d}(\mathbb{R}^{n}))$.
By Proposition 2.3, for every $\lambda>0$ there exists an $\lambda$-Lipschitz map $f_{\lambda}:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ with
$\lambda^{p}|\{f\neq f_{\lambda}\}|\leq C\int_{\{f\neq f_{\lambda}\}}(|Df|^{k}+|f|^{k}).$
We have that $Df=Df_{\lambda}$ almost everywhere outside $\{f\neq f_{\lambda}\}$, cf. Proposition 2.3. Consequently
$\Big|\int\beta\wedge(f_{\lambda}^{*}\omega-f^{*}\omega)\Big|$ $\leq C\|\beta\|_{\infty}\int_{\{f\neq f_{\lambda}\}}(\|f_{\lambda}^{*}\omega\|+\|f^{*}\omega\|)$
$\leq C\|\beta\|_{\infty}\|\omega\|_{\infty}\int_{\{f\neq f_{\lambda}\}}(\lambda^{k}+|Df|^{k})\mathrm{d}x$
$\leq C\|\beta\|_{\infty}\|\omega\|_{\infty}\int_{\{f\neq f_{\lambda}\}}(|Df|^{k}+|f|^{k})\stackrel{{\scriptstyle\lambda\to\infty}}{{\longrightarrow}}0.$
for all $\beta\in\Omega_{c}^{m-k}(U)$. Similarly (either since $k+1\leq p$ or since $\mathrm{d}\omega=0$) we obtain
$\Big|\int_{U}\alpha\wedge(f_{\lambda}^{*}(\mathrm{d}\omega)-f^{*}(\mathrm{d}\omega))\Big|\stackrel{{\scriptstyle\lambda\to\infty}}{{\longrightarrow}}0.$
for all $\alpha\in\Omega_{c}^{m-k-1}(U)$. These identities together with Theorem 4.7 imply
$\int_{U}\mathrm{d}\alpha\wedge f^{*}\omega$ $=\lim_{\lambda\to\infty}\int_{U}\mathrm{d}\alpha\wedge f_{\lambda}^{*}\omega=(-1)^{k+1}\lim_{\lambda\to\infty}\int_{U}\alpha\wedge f_{\lambda}^{*}(\mathrm{d}\omega)$
$=(-1)^{k+1}\int_{U}\alpha\wedge f^{*}(\mathrm{d}\omega)$
for all $\alpha\in\Omega_{c}^{m-k-1}(U)$. ∎
5. Multivalued inverse of QR map as a QR curve
Throughout this section, $\omega=\operatorname{tr}(\operatorname{vol}_{\mathbb{R}^{n}})$ denotes the canonical $n$-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$.
Lemma 5.1.
Let $f\in W^{1,n}_{loc}(\Omega,\mathcal{A}_{d}(\mathbb{R}^{n}))$, where $\Omega$ is an open set. Then
$\star f^{*}\omega=\sum_{j}^{d}\det(Df_{j})\quad\textrm{ a.e. on }\Omega.$ (5.1)
where $Df=\llbracket Df_{1},\ldots,Df_{d}\rrbracket$ is the weak differential of $f$. In particular, we have the inequality
$|\star f^{*}\omega|\leq\mathbf{J}f\quad\textrm{a.e. on }U.$ (5.2)
Proof.
Let $Df=\llbracket Df_{1}(x),\ldots,Df_{d}(x)\rrbracket$ be the differential of $f$ at $x$. By Definitions 4.4 and 4.6 we have that
$(f^{*}\omega)_{x}(v_{1},\ldots,v_{n})$ $=\omega_{f(x)}(Df(x)v_{1},\ldots,Df(x)v_{n})$
$=\sum_{j}^{d}Df_{j}(x)v_{1}\wedge\cdots\wedge Df_{j}(x)v_{n}$
$=\sum_{j}^{d}\det(Df_{j}(x))v_{1}\wedge\cdots\wedge v_{n},\quad v_{1},\ldots,v_{n}\in T_{x}U,$
which implies (5.1). The estimate (5.2) follows from (5.1) together with Lemma 2.4 and Minkowski’s determinant inequality:
$|\star f^{*}\omega|$ $\leq\sum_{j}^{d}|\det(Df_{j})|=\sum_{j}^{d}[\det(Df_{j}^{T}Df_{j})]^{1/2}$
$\leq\Big(\sum_{j}^{d}[\det(Df_{j}^{T}Df_{j})]^{1/n}\Big)^{n/2}\leq\det\Big(\sum_{j}^{d}Df_{j}^{T}Df_{j}\Big)^{1/n\ \cdot\ n/2}=\mathbf{J}f.$
Proof of Theorem 1.2.
The fact that $\accentset{\leftarrow}{f}\in W^{1,n}_{loc}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n})$ follows from Theorem 3.7 and Corollary 3.8. For every $y\in f(\Omega)\setminus f(B_{f})$ we have $f^{-1}(y)=d$ and we can find $r>0$ such that $f^{-1}:B(y,r)\to U_{f}(x,r)$ is a homeomorphism for each $x\in f^{-1}(y)$. Choose a numbering $f^{-1}(y)=\{x_{1},\ldots,x_{d}\}$ and denote $U_{j}=U_{f}(x_{j},r)$ and $g_{j}=f^{-1}:B(y,r)\to U_{j}$. Then
$\accentset{\leftarrow}{f}|_{B(y,r)}=\llbracket g_{1},\ldots,g_{d}\rrbracket,$
and each $g_{j}$ is a $K_{I}(f)$-quasiconformal map between domains in $\mathbb{R}^{n}$, in particular
$\|Dg_{j}\|^{n}\leq K_{I}J_{g_{j}}=K_{I}\star g_{j}^{*}(\mathrm{d}x).$
Observe that $\star\accentset{\leftarrow}{f}^{*}\omega=\sum_{j}^{d}J_{g_{j}}$ and $|D\accentset{\leftarrow}{f}|^{2}\leq\sum_{j}^{d}\|Dg_{j}\|^{2}$ a.e. on $B(y,r)$. Thus
$|D\accentset{\leftarrow}{f}|^{n}$ $\leq\Big(\sum_{j}^{d}\|Dg_{j}\|^{2}\Big)^{n/2}\leq d^{n/2-1}\sum_{j}^{d}\|Dg_{j}\|^{n}$
$\leq K_{I}d^{n/2-1}\sum_{j}^{d}J_{g_{j}}=d^{n/2-1}K_{I}\star\accentset{\leftarrow}{f}^{*}\omega$
a.e. in $B(y,r)$. Since $|f(B_{f})|=0$ by [26, Chapter II, Proposition 4.14], the inequality holds a.e. on $f(\Omega)$ as claimed. ∎
Remark 5.2.
An inspection of the proof yields the following inequality:
$\mathbf{J}\accentset{\leftarrow}{f}\leq K_{I}\star\accentset{\leftarrow}{f}^{*}\omega\quad\textrm{ a.e. on }f(\Omega).$ (5.3)
This a reverse version (up to the multiplicative constant $K_{I}$) of the inequality in Lemma 5.1. Unlike Lemma 5.1, (5.3) does not hold for general Sobolev maps.
The following higher integrability result for $\accentset{\leftarrow}{f}$ implies part of Corollary 1.3.
Theorem 5.3.
If $f:\Omega\to\mathbb{R}^{n}$ is a non-constant quasiregular map of degree $d$, then there exists $p>n$ such that $\accentset{\leftarrow}{f}\in W_{loc}^{1,p}(f(\Omega),\mathcal{A}_{d}(\Omega))$.
Proof.
The proof is identical to Lemma 6.1, Lemma 6.2 and Proposition 6.3 in [21]. The only place a minor argument is needed is at the beginning of Lemma 6.1, where it is stated that $\omega=\mathrm{d}\tau$ for a suitable (n1)(n-1)-form $\tau$. We need to ensure that $\tau$ can be chosen $S_{d}$-invariant. But this follows from Lemma 4.2: $\omega=P_{S_{d}}\omega=\mathrm{d}(P_{S_{d}}\tau)$, and $\tau^{\prime}=P_{S_{d}}\tau$ is $S_{d}$-invariant. The rest of the proof(s) proceed exactly as in [21]. Note that $\mathrm{d}(f^{*}\tau^{\prime})=f^{*}\omega$ (a fact needed in the proof in [21]) holds by Theorem 1.1. ∎
The following theorem states that the multi-valued inverse of a QR map is a quasiminimizer of the $n$-energy and the parametrized volume, respectively, which are defined for $u\in W^{1,n}_{loc}(\Omega,\mathcal{A}_{d}(\mathbb{R}^{n}))$ and $V\subset\Omega$ as follows.
$E_{n}(u;V):=\int_{V}\|Du\|^{n}\mathrm{d}x,\quad A(u;V):=\int_{V}\mathbf{J}u\ \mathrm{d}x.$
Theorem 5.4.
Let $\Omega\subset\mathbb{R}^{n}$ be a domain and $f:\Omega\to\mathbb{R}^{n}$ a proper QR map. If $u\in W_{loc}^{1,n}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ is such that $u=\accentset{\leftarrow}{f}$ a.e. outside a compact set $K\subset\Omega$, then
$E_{n}(\accentset{\leftarrow}{f};K)$ $\leq d^{n/2-1}K_{I}E_{n}(u;K)$
$A(\accentset{\leftarrow}{f};K)$ $\leq K_{I}A(u;K)$
Recall that $K_{I}=K_{I}(f)$ is the inner distortion of $f$.
Proof.
Let $\tau$ be an (n1)(n-1)-form on $\mathcal{A}_{d}(\mathbb{R}^{n})$ such that $\omega=\mathrm{d}\tau$. Since $u=\accentset{\leftarrow}{f}$ a.e. on $\Omega\setminus K$, we have that $u^{*}\omega=\accentset{\leftarrow}{f}^{*}\omega$ and $u^{*}\tau=\accentset{\leftarrow}{f}^{*}\tau$ a.e. on $\Omega\setminus K$.
Set $K_{\varepsilon}:=\{x:\ \operatorname{dist}(K,x)\leq\varepsilon\}$, and note that $K_{\varepsilon}$ is a finite perimeter set for a.e. $\varepsilon$. Moreover $u^{*}\tau=\accentset{\leftarrow}{f}^{*}\tau$ a.e. on $\partial^{*}K_{\varepsilon}$ for a.e. $\varepsilon$ by a Fubini-type argument. Thus, by a generalized Stoke’s theorem, we have
$\int_{K_{\varepsilon}}u^{*}\omega=\int_{K_{\varepsilon}}\mathrm{d}(u^{*}\tau)=\int_{\partial^{*}K_{\varepsilon}}u^{*}\tau=\int_{\partial^{*}K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\tau=\int_{K_{\varepsilon}}\mathrm{d}(\accentset{\leftarrow}{f}^{*}\tau)=\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega$ (5.4)
for a.e. $\varepsilon>0$. (Note that we use Theorem 1.1 in the first and last equalities above.) Indeed, $\llbracket K_{\varepsilon}\rrbracket$ is a normal $n$-current on $\mathbb{R}^{n}$ and $\partial\llbracket K_{\varepsilon}\rrbracket=\llbracket\partial^{*}K_{\varepsilon}\rrbracket$ whenever $K_{\varepsilon}$ has finite perimeter, and in this case we have by [15, Theorem 1.1] the following: for each $v\in W^{1,n}(f(\Omega),\mathcal{A}_{d}(\mathbb{R}^{n}))$ the identity $\partial(v_{\ast}\llbracket K_{\varepsilon}\rrbracket)=v_{\ast}(\partial\llbracket K_{\varepsilon}\rrbracket)=v_{\ast}(\llbracket\partial^{*}K_{\varepsilon}\rrbracket)$ holds a.e. $\varepsilon$. This implies
$\int_{K_{\varepsilon}}\mathrm{d}(v^{*}\beta)=\int_{\partial^{*}K_{\varepsilon}}v^{*}\beta$
for smooth (n1)(n-1)-forms $\beta$ on $\mathcal{A}_{d}(\mathbb{R}^{n})$ (cf. also [15, Lemma 5.3]), which in turn implies (5.4) for a.e. $\varepsilon$. Using (5.4) we may estimate
$E_{n}(\accentset{\leftarrow}{f};K)$ $\leq\int_{K_{\varepsilon}}\|D\accentset{\leftarrow}{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K_{\varepsilon}}\star\accentset{\leftarrow}{f}^{*}\omega=d^{n/2-1}K_{I}\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega$
$=d^{n/2-1}K_{I}\int_{K_{\varepsilon}}u^{*}\omega\leq d^{n/2-1}K_{I}E_{n}(u;K_{\varepsilon})$
for a.e. $\varepsilon>0$. The last inequality above follows from the pointwise inequalities $\star u^{*}\omega\leq\mathbf{J}u\leq\|Du\|^{n}$, cf. Lemma 5.1. Since $\mathcal{L}^{n}(K_{\varepsilon}\setminus K)\to 0$ as $\varepsilon\to 0$, we obtain
$E_{n}(\accentset{\leftarrow}{f};K)\leq d^{n/2-1}K_{I}E_{n}(u;K).$
Similarly,
$A(\accentset{\leftarrow}{f};K)\leq\int_{K_{\varepsilon}}\mathbf{J}\accentset{\leftarrow}{f}\mathrm{d}x\leq K_{I}\int_{K_{\varepsilon}}\accentset{\leftarrow}{f}^{*}\omega=K_{I}\int_{K_{\varepsilon}}u^{*}\omega\leq K_{I}A(u;K_{\varepsilon})$
(cf. Remark 5.2 and Lemma 5.1) yielding
$A(\accentset{\leftarrow}{f};K)\leq K_{I}A(u;K)$
at the limit $\varepsilon\to 0$. This completes the proof. ∎
Remark 5.5.
The use of pull-back theory and the generalized Stokes theorem could also be used to give another proof of Theorem 5.3. Indeed, it yields a reverse Hölder inequality without a Caccioppoli inequality, and the higher integrability is then a standard consequence of a theorem of Gehring. We omit the details.
We end this section with a proof of Corollary 1.3, which is essentially a consequence of Theorems 5.3 and 5.4.
Proof of Corollary 1.3.
Claim (i) follows immediately from Theorem 5.3, Corollary A.2(i) and the observation that $g_{f}=b\circ\accentset{\leftarrow}{f}$.
To prove (ii), let $h\in W^{1,n}_{loc}(f(\Omega),\mathbb{R}^{n})$ be such that $\{h\neq g_{f}\}$ is essentially contained in a compact subset $K\subset f(\Omega)$. A straightforward calculation shows that if $R>0$, then the multivalued map $R\accentset{\leftarrow}{f}$, defined by $R\accentset{\leftarrow}{f}(x)=\sum_{y\in f^{-1}(x)}\iota(f,x)\llbracket Rx\rrbracket$, is an energy quasi-minimizer with the same constant as $\accentset{\leftarrow}{f}$ and $\|D(R\accentset{\leftarrow}{f})\|=R\|D\accentset{\leftarrow}{f}\|$. Now consider the multivalued function
$h_{R}=\mathcal{L}^{-1}\circ(R\cdot h,\pi_{\mathcal{Z}}\circ\accentset{\leftarrow}{f})$
Then $\{h_{R}\neq\accentset{\leftarrow}{f}\}$ is essentially contained in $K$, and by Theorem 5.4 and the discussion above we have
$\int_{K}R^{n}\|D\accentset{\leftarrow}{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh_{R}\|^{n}\mathrm{d}x.$
By Corollary A.2(ii) this implies
$\int_{K}d^{n/2}R^{n}\|Dg_{f}\|^{n}\leq d^{n/2-1}K_{I}\int_{K}(dR^{2}\|Dh\|^{2}+\|D(\pi_{\mathcal{Z}}\circ\accentset{\leftarrow}{f})\|^{2})^{n/2}\mathrm{d}x.$
Dividing by $R^{n}$ and taking the limit $R\to+\infty$ we obtain
$\int_{K}\|Dg_{f}\|^{n}\mathrm{d}x\leq d^{n/2-1}K_{I}\int_{K}\|Dh\|^{n}\mathrm{d}x.$
6. Multivalued inverse of QR map as a QC map
Throughout this section we assume that $f:\Omega\to\mathbb{R}^{n}$ is a proper quasiregular map of degree $d$. The main result of this section is the geometric quasiconformality of the multi-valued inverse onto its image, Theorem 1.4, whose proof is presented in Section 6.2.
6.1. Image of multivalued inverse
Consider the space $\Omega_{f}:=\accentset{\leftarrow}{f}(f(\Omega))$ equipped with the metric $d_{\mathcal{A}}$ from $\mathcal{A}_{d}(\Omega)$. We moreover equip $\Omega_{f}$ with its Hausdorff $n$-measure $\mathcal{H}^{n}$. Since $\accentset{\leftarrow}{f}:f(\Omega)\to\Omega_{f}$ is a homeomorphism, $\Omega_{f}$ is an open $n$-manifold. Moreover, $\mathcal{H}^{n}$ is locally finite by the area-formula.
Lemma 6.1.
Let $\mathcal{H}^{n}$ be the Hausdorff $n$-measure on $\Omega_{f}$. Then
  • (i)
    $\Omega_{f}$ is $n$-rectifiable;
  • (ii)
    $\mathcal{H}^{n}$ is locally finite;
  • (iii)
    $\mathcal{H}^{n}(\accentset{\leftarrow}{f}(f(B_{f})))=0$;
Proof.
All three claims follow from Proposition 3.9. ∎
Next, we establish the upper Ahlfors $n$-regularity of the Hausdorff measure on $\Omega_{f}$.
Proposition 6.2.
For any compact set $K\subset\Omega_{f}$ there exists $r_{K}>0$ such that
$\mathcal{H}^{n}(B_{\Omega_{f}}(z,r))\leq\omega_{n}d^{n/2}K_{I}K_{O}r^{n},\quad z\in K,\ r<r_{K}.$
Proof.
For any $z\in K$, let $r_{z}>0$ be such that
$(\accentset{\leftarrow}{f}\circ f)^{-1}B(z,r)=\bigcup_{x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z)}U_{\accentset{\leftarrow}{f}\circ f}(x,r)$
and $U_{\accentset{\leftarrow}{f}\circ f}(x,r)$ is a normal neighbourhood of $x$ for each $x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z)$. Let $r_{K}$ be the Lebesgue number of the cover $\{B(z,r_{z})\}_{z\in K}$. For any $z\in K$ there exists $z_{0}\in K$ such that $B(z,r)\subset B(z_{0},r_{z_{0}})$ for $r<r_{K}$. Note that
$U_{\accentset{\leftarrow}{f}\circ f}(x,r)\subset B(x,r),\quad x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z).$
Indeed, for any $y\in U_{F\circ f}(x,r)$ we have
$|y-x|\leq d(\accentset{\leftarrow}{f}(f(y)),F\accentset{\leftarrow}{f}(f(x))=d(\accentset{\leftarrow}{f}\circ f(y),z)<r.$
Now the area formula (3.6) yields
$\mathcal{H}^{n}(B(z,r))$ $=\int_{\accentset{\leftarrow}{f}^{-1}B(z,r)}\mathbf{J}\accentset{\leftarrow}{f}\mathrm{d}y\leq K_{I}K_{O}\int_{\accentset{\leftarrow}{f}^{-1}B(z,r)}(f_{\ast}(|Df|^{-2}))^{n/2}\mathrm{d}y$
$\leq d^{n/2-1}K_{I}K_{O}\int_{\accentset{\leftarrow}{f}^{-1}B(z,r)}f_{\ast}(|Df|^{-n})\mathrm{d}y$
$=d^{n/2-1}K_{I}K_{O}\int_{f^{-1}(\accentset{\leftarrow}{f}^{-1}B(z,r))}|Df|^{-n}\mathbf{J}f\mathrm{d}x$
$=d^{n/2-1}K_{I}K_{O}\sum_{x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z)}\int_{U_{\accentset{\leftarrow}{f}\circ f}(x,r)}\frac{\mathbf{J}\accentset{\leftarrow}{f}}{|Df|^{n}}\mathrm{d}x$
$\leq d^{n/2-1}K_{I}K_{O}\sum_{x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z)}|B(x,r)|\leq\omega_{n}d^{n/2}K_{I}K_{O}r^{n}.$
Remark 6.3.
Using Proposition 3.4 (or arguing along the lines in the proof above of Proposition 6.2) we can obtain the lower bound
$\mathcal{H}^{n}(B(z,r))\geq\frac{1}{K_{O}d}\sum_{x\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z)}|U_{\accentset{\leftarrow}{f}\circ f}(x,r)|,\quad z\in K,\ r<r_{K}.$
A lower bound for $\frac{|U_{\accentset{\leftarrow}{f}\circ f}(x,r)|}{r^{n}}$, together with the LLC-property, would imply a local $(1,1)$-Poincaré inequality for $\Omega_{f}$ by a deep result of Semmes [27].
In the absence of the lower bound we instead establish, in the next proposition, an infinitesimal $n$-Poincaré inequality for $\Omega_{f}$. This is a weaker qualitative notion compared to $p$-Poincaré inequalities, but it is sufficient in order to obtain the geometric quasiconformality of the multi-valued inverse.
A metric measure space $X=(X,d,\mu)$ is said to satisfy an infinitesimal $p$-Poincaré inequality ($p$-PI for short), if
$|Dg|_{p}=\operatorname{Lip}g\quad\mu\textrm{-a.e. on }X$ (6.1)
for every $g\in\operatorname{LIP}(X)$. Here $|Dg|_{p}$ denotes the minimal $p$-weak upper gradient of $g$.
Proposition 6.4.
The space $\Omega_{f}=(\Omega_{f},d_{\mathcal{A}},\mathcal{H}^{n})$ satisfies (6.1) with $p=n$ for every $h\in\operatorname{LIP}(\Omega_{f})$.
In particular, $\Omega_{f}$ is an $n$-dimensional differentiability space, cf. [14, Theorem 1.1]. In the proof we use the terminology and techniques from [3]. Since these are used only in this proof we do not recall them here but refer the reader instead to [3].
Proof.
The map $F=\accentset{\leftarrow}{f}\circ f:\Omega\to\Omega_{f}$ is a proper branched cover of degree $d$. It is almost everywhere metrically differentiable and satisfies Lusin’s properies $(N)$ and $(N^{-1})$. Moreover $F$ also satisfies the modulus inequality
$\operatorname{Mod}_{n}\Gamma\leq C\operatorname{Mod}_{n}F(\Gamma)$
for path families in $\Omega$, cf. Proposition 6.6 and [12, Lemma 3.6].
By Lemma 6.1 we have that $\mathcal{H}^{n}(F(B_{f}))=0$. For every $z_{0}\in\Omega_{f}\setminus F(B_{f})$ there is $r>0$ so that $F_{0}=F|_{U_{F}(x_{0},r)}:U_{F}(x_{0},r)\to B(z_{0},r)=:B_{0}$ is an analytically quasiconformal homeomorphism for every $x_{0}\in F^{-1}(z_{0})$. Since $d_{\mathcal{A}}(F_{x}(y),F_{x}(y^{\prime}))\geq|y-y^{\prime}|$ we obtain that $F_{0}^{-1}:B_{0}\to U_{0}:=U_{F}(x_{0},r)$ is 1-Lipschitz.
We claim that $(B_{0},F_{0}^{-1})$ is both a Cheeger chart and an $n$-weak chart. Since $F_{0}$ is metrically differentiable everywhere and has $(N)$ and $(N^{-1})$, it is not difficult to see that $(B_{0},F_{0})$ is a Cheeger chart. [^2] Indeed, a direct calculation yields that $(B_{0},F_{0}^{-1})$ is a weak Cheeger chart. Since $B_{0}$ is open, it follows that $(B_{0},F_{0}^{-1})$ is a Cheeger chart. See [12] for the terminology. It also follows that, if $(B_{0},F_{0})$ is $n$-independent (in the sense of [3]) then it is maximal, and thus an $n$-weak chart.
Observe that, for $\operatorname{Mod}_{n}$-a.e. $\gamma$ in $\Omega$ we have
$|(F_{0}\circ\gamma)_{t}^{\prime}|=\operatorname{\operatorname{md}}_{\gamma_{t}}F_{0}(\gamma_{t}^{\prime})\quad\textrm{a.e. }t.$ (6.2)
Let $\Gamma_{0}$ be the $\operatorname{Mod}_{n}$-null path family where (6.2) fails. Given $\xi\in(\mathbb{R}^{n})^{*}$ with norm 1 and $v\in S^{n-1}$ such $\xi(v)=1$, let $\Gamma_{v}$ be the path family of curves in $U_{0}$ with $\gamma_{t}^{\prime}=v$. Then $0<\operatorname{Mod}_{n}(\Gamma_{v}\setminus\Gamma_{0})\leq C\operatorname{Mod}_{n}F(\Gamma_{v}\setminus\Gamma_{0})$ and, for $\operatorname{Mod}_{n}$-a.e. curve $\gamma$ in $F(\Gamma_{v}\setminus\Gamma_{0})$, we have
$|\gamma_{t}^{\prime}|=\operatorname{\operatorname{md}}_{F_{0}^{-1}(\gamma_{t})}F_{0}(v),\quad(F_{0}^{-1}\circ\gamma)_{t}^{\prime}=v\quad\textrm{a.e. }t.$
Thus
$\frac{\xi((F_{0}^{-1}\circ\gamma)_{t}^{\prime})}{|\gamma_{t}^{\prime}|}=\frac{1}{\operatorname{\operatorname{md}}_{F_{0}^{-1}(\gamma_{t})}F_{0}(v)}\geq\frac{1}{|DF_{0}|(F_{0}^{-1}(\gamma_{t}))}.$
This implies
$|D(\xi\circ F_{0}^{-1})|_{n}(x)\geq\frac{1}{|DF_{0}|(F_{0}^{-1}(x))}\quad\textrm{a.e. }x\in B_{0}.$
Consequently $F_{0}^{-1}$ is $n$-independent in $B_{0}$, completing the proof that $(B_{0},F_{0}^{-1})$ is a $n$-weak chart.
Now since $(B_{0},F_{0}^{-1})$ is an $n$-dimensional Cheeger chart and $n$-weak chart, it follows from [3, Theorem 1.11] and [14, Theorem 1.3] that $\operatorname{Lip}g=|Dg|_{n}$ $\mathcal{H}^{n}$-a.e. on $\Omega_{f}$ for any $g\in\operatorname{LIP}(\Omega_{f})$. This completes the proof. ∎
6.2. Quasiconformality of multi-valued inverse
It is not difficult to show that $\accentset{\leftarrow}{f}$ is metrically quasiconformal, i.e. $H(\accentset{\leftarrow}{f},x)\leq H$ for all $x\in f(\Omega)$, where $H(\accentset{\leftarrow}{f},x)=\limsup_{r\to 0}\frac{L_{\accentset{\leftarrow}{f}}(x,r)}{l_{\accentset{\leftarrow}{f}}(x,r)}$ and
$L_{\accentset{\leftarrow}{f}}(x,r)$ $=\sup\{d_{\mathcal{A}}(\accentset{\leftarrow}{f}(x),\accentset{\leftarrow}{f}(y)):\ d(x,y)\leq r\}$
$l_{\accentset{\leftarrow}{f}}(x,r)$ $=\sup\{d_{\mathcal{A}}(\accentset{\leftarrow}{f}(x),\accentset{\leftarrow}{f}(y)):\ d(x,y)\geq r\}.$
In fact we have the following.
Proposition 6.5.
The multi-valued inverse $\accentset{\leftarrow}{f}$ of $f$ satisfies, for small $r>0$:
$H_{\accentset{\leftarrow}{f}}(y,r)^{2}\leq\sum_{x\in f^{-1}(y)}H_{f}^{*}(x,r)^{2},\quad y\in f(\Omega).$
Here $H^{*}(f,x)=\limsup_{r\to 0}\frac{L_{f}^{*}(x,r)}{l_{f}^{*}(x,r)}$ and
$L_{f}^{*}(x,r)$ $=\sup\{d_{\mathcal{A}}(\accentset{\leftarrow}{f}(x),\accentset{\leftarrow}{f}(y)):\ y\in U_{f}(x,r)\}$
$l_{f}^{*}(x,r)$ $=\sup\{d_{\mathcal{A}}(\accentset{\leftarrow}{f}(x),\accentset{\leftarrow}{f}(y)):\ y\in\partial U_{f}(x,r)\}.$
Proof.
Denote $f^{-1}(y)=\{x_{1},\ldots,x_{m}\}$ where $\iota(f,x_{1})+\cdots+\iota(f,x_{m})=d$. Let $r$ be small enough so that $f^{-1}B(y,r)=\bigcup_{j}^{m}U_{f}(x_{j},r)$ and $U_{j}:=U_{f}(x_{j},r)$ is a normal domain of $x_{j}$, $j=1,\ldots,m$. For $z\in B(y,r)$ we have $f^{-1}(z)=\bigcup_{j}^{m}f^{-1}(z)\cap U_{j}$ and
$d_{\mathcal{A}}^{2}(\accentset{\leftarrow}{f}(y),\accentset{\leftarrow}{f}(z))=\sum_{j}^{m}\sum_{x^{\prime}\in f^{-1}(z)\cap U_{j}}|x_{j}-x^{\prime}|^{2}\leq\sum_{j}^{m}\iota(f,x_{j})L^{*}_{f}(x_{j},r)^{2}.$
This implies $L_{\accentset{\leftarrow}{f}}(y,r)^{2}\leq f_{\ast}L^{*}_{f}(\cdot,r)(y)$. Similarly, for $z\in\partial B(y,r)$,
$d_{\mathcal{A}}^{2}(\accentset{\leftarrow}{f}(y),\accentset{\leftarrow}{f}(z))=\sum_{j}^{m}\sum_{x^{\prime}\in f^{-1}(z)\cap U_{j}}|x_{j}-x^{\prime}|^{2}\geq\sum_{j}^{m}\iota(f,x_{j})l^{*}_{f}(x_{j},r)^{2},$
implying $l_{\accentset{\leftarrow}{f}}(y,r)^{2}\geq f_{\ast}(l^{*}_{f}(\cdot,r)^{2})$. Together these estimates yield the claim. ∎
The next proposition states that $\accentset{\leftarrow}{f}$ is analytically quasiconformal.
Proposition 6.6.
The multi-valued inverse $\accentset{\leftarrow}{f}:f(\Omega)\to\Omega_{f}$ of $f$ is an analytically quasiconformal homeomorphism, i.e.
$\operatorname{Mod}_{n}\Gamma\leq K_{I}K_{O}\operatorname{Mod}_{n}\accentset{\leftarrow}{f}(\Gamma)$
for every path family $\Gamma$ in $f(\Omega)$. Moreover
$(f_{\ast}(|Df|^{-2}))^{n/2}\leq\mathbf{J}\accentset{\leftarrow}{f}\leq{K_{I}K_{O}}(f_{\ast}(|Df|^{-2}))^{n/2}\quad\textrm{almost everywhere.}$ (6.3)
Proof.
Denote $H(y):=(f_{\ast}(|Df|^{-2}))^{1/2}$ and observe that
$\lim_{h\to 0}\frac{d_{\mathcal{A}}(\accentset{\leftarrow}{f}(y+hv),\accentset{\leftarrow}{f}(y))}{|h|}=\operatorname{\operatorname{md}}_{y}\accentset{\leftarrow}{f}(v)\quad\textrm{ a.e. }y\in f(\Omega)$
for each $v\in\mathbb{S}^{n-1}$. By considering the family $\Gamma_{v}=\{\gamma\subset f(\Omega):\gamma^{\prime}=v\}$, it follows from a Fubini type argument and Theorem 3.7 that
$H(y)\leq\operatorname{\operatorname{md}}_{y}\accentset{\leftarrow}{f}(v)\leq(K_{I}K_{O})^{1/n}H(y)\quad\textrm{ for all }v\in\mathbb{S}^{n-1}.$ (6.4)
This implies (6.3). The remaining claim follows from (6.4). Let $\Gamma$ be a path family in $f(\Omega)$ and let $\rho$ be admissible for $\accentset{\leftarrow}{f}(\Gamma)$. If $\Gamma_{0}$ is the exceptional family in Theorem 3.7 then $\operatorname{Mod}_{v}\Gamma_{0}=0$ and, for any $\gamma\in\Gamma\setminus\Gamma_{1}$ we have
$1\leq\int_{\accentset{\leftarrow}{f}\circ\gamma}\rho\mathrm{d}s$ $=\int_{0}^{1}\rho(\accentset{\leftarrow}{f}(\gamma_{t}))|(\accentset{\leftarrow}{f}\circ\gamma)_{t}^{\prime}|\mathrm{d}t\leq(K_{I}K_{O})^{1/n}\int_{0}^{1}\rho\circ\accentset{\leftarrow}{f}(\gamma_{t})H(\gamma_{t})|\gamma_{t}^{\prime}|\mathrm{d}t$
$=(K_{I}K_{O})^{1/n}\int_{\gamma}(\rho\circ\accentset{\leftarrow}{f})\cdot H\mathrm{d}s.$
Thus $(K_{I}K_{O})^{1/n}(\rho\circ\accentset{\leftarrow}{f})H$ is admissible for $\Gamma\setminus\Gamma_{0}$ implying
$\operatorname{Mod}_{n}\Gamma$ $=\operatorname{Mod}_{n}\Gamma\setminus\Gamma_{0}\leq K_{I}K_{O}\int_{f(\Omega)}\rho^{n}\circ\accentset{\leftarrow}{f}H^{n}\mathrm{d}y$
$\leq K_{I}K_{O}\int_{f(\Omega)}\rho^{n}\circ\accentset{\leftarrow}{f}\mathbf{J}\accentset{\leftarrow}{f}\mathrm{d}y=K_{I}K_{O}\int_{\Omega_{f}}\rho^{n}\mathrm{d}\mathcal{H}^{n}.$
Taking infimum over $\rho$ yields $\operatorname{Mod}_{n}\Gamma\leq K_{I}K_{O}\operatorname{Mod}_{n}\accentset{\leftarrow}{f}(\Gamma)$ and completes the proof. ∎
The other inequality requires more work. It is based on the following proposition and an induction argument involving a similar approximation as in the proof of Theorem 4.7.
Proposition 6.7.
Let $h:X\to Y$ be a homeomorphism in $N^{1,n}_{loc}(X,Y)$ with $|Dh|^{n}\leq K\ \mathbf{J}h$ and Lusin’s property $(N)$ between $n$-rectifiable spaces satisfying the infinitesimal $n$-PI. If $h^{-1}\in N^{1,n}_{loc}(Y,X)$ and $h^{-1}$ has Lusin’s property $(N)$, then $|D(h^{-1})|^{n}\leq K\mathbf{J}(h^{-1})$.
Proof.
For a Sobolev map $f\in N^{1,n}(Z,W)$ with Lusin’s properties $(N)$ and $(N^{-1})$ between $n$-rectifiable spaces satisfying the infinitesimal $n$-PI, the minimal upper gradient $|Df|$ of $f$ coincides $\mathcal{H}^{n}$-a.e. $x\in Z$ with $\|D_{x}f\|_{op}$, where $D_{x}f$ is the approximate differential of $f$ at $x$, and the volume Jacobian of [30] coincides with $\mathbf{J}f=\mathbf{J}(D_{x}f)$. Thus the inequality $|Df|^{n}\leq K\ \mathbf{J}f$ is equivalent with the modulus inequality
$\operatorname{Mod}_{n}\Gamma\leq K\operatorname{Mod}_{n}f(\Gamma)$
for path families $\Gamma$ in $X$.
Now, by rectifiability and the infinitesimal Poincaré inequality we have that Gigli’s co-tangent module $L^{n}(T^{*}Z)$ coincides with the $n$-integrable sections $\Gamma_{n}(T^{*}Z)$ of the co-tangent bundle given by the rectifiable atlas, cf. [14, Theorems 1.2 and 1.3], $Z=X,Y$. The claim now follows from [12, Theorem 3.13]. Indeed, Using the above identification (and the fact that $h$ has Lusin’s properties $(N)$ and $(N^{-1})$ we have that the pull-back operator $h^{*}$ constructed in [12, Theorem 3.4] coincides with $(Dh)^{*}:T^{*}Y\to T^{*}X$, where $Dh:TX\to TY$ is the bundle map given by the a.e. defined approximate differential of $h$. Since
$\|Dh\|_{op}\leq K\mathbf{J}(Dh),$
$(Dh)^{*}$ is an isomorphism, proving condition (i) in [12, Theorem 3.13], while condition (ii) in [12, Theorem 3.13] is satisfied since $h^{-1}\in N_{loc}^{1,p}(Y,X)$. ∎
Theorem 6.8.
If $f:\Omega\to\mathbb{R}^{n}$ is a quasiregular map of finite degree $d>0$, then the multi-valued inverse satisfies $\accentset{\leftarrow}{f}^{-1}\in N_{loc}^{1,n}(\Omega_{f},\mathbb{R}^{n})$ and
$|D\accentset{\leftarrow}{f}^{-1}|\leq\frac{(K_{I}K_{O})^{1/n}}{H\circ\accentset{\leftarrow}{f}^{-1}},\quad|D\accentset{\leftarrow}{f}^{-1}|^{n}\leq K_{I}K_{O}\mathbf{J}(\accentset{\leftarrow}{f}^{-1}).$
Here
$H(y):=(f_{\ast}(|Df|^{-2})(y))^{1/2},\quad y\in f(\Omega).$
Proof.
We will prove the claim by induction on the degree of $f$. Denote by $I(k)$ the induction hypothesis: $\accentset{\leftarrow}{f}^{-1}\in N^{1,n}_{loc}(U_{f},\mathbb{R}^{n})$ and $|D\accentset{\leftarrow}{f}^{-1}|\leq\frac{C(f)}{H\circ\accentset{\leftarrow}{f}^{-1}}$ for every quasiregular map $f:U\to\mathbb{R}^{n}$ of degree $\leq k$ from an open set $U\subset\mathbb{R}^{n}$, where $C(f)=(K_{I}(f)K_{O}(f))^{1/n}$. When $k=1$, the claim is classical. Indeed, a locally homeomorphic quasiregular map is quasiconformal [26, Chapter III, Corollary 3.8], and its inverse is quasiconformal with the correct bound [26, Chapter II, Corollary 6.5]. Thus $I(1)$ holds.
Suppose now that $I(k)$ holds for some $k\in\mathbb{N}$. We will prove that $I(k+1)$ holds. Let $d=k+1$ and let $f:\Omega\to\mathbb{R}^{n}$ be a quasiregular map of degree $d$. Set $A=\Omega_{f}\cap\operatorname{diag}\mathcal{A}_{d}(\Omega)$ and note that, for all $z_{0}\in\Omega_{f}\setminus A$ there exists $x_{0}\in(\accentset{\leftarrow}{f}\circ f)^{-1}(z_{0})$ and $r_{0}>$ such that $\iota(f,x_{0})<d$ and the map
$f_{0}:=f|_{U_{\accentset{\leftarrow}{f}\circ f}(x_{0},r_{0})}:U_{\accentset{\leftarrow}{f}\circ f}(x_{0},r_{0})\to F^{-1}B(z_{0},r_{0})$
is a quasiregular map of degree $\leq k$. Denote $U_{0}=U_{\accentset{\leftarrow}{f}\circ f}(x_{0},r_{0})$, $B_{0}=B(z_{0},r_{0})$, $\accentset{\leftarrow}{f}_{0}:F^{-1}B_{0}\to\Omega_{f_{0}}$ (the multivalued inverse of $f_{0}$), and $H_{0}=\big(f_{0\ast}(|Df_{0}|^{-2}))^{1/2}$. By induction hypothesis $\accentset{\leftarrow}{f}_{0}^{-1}\in N^{1,n}_{loc}((U_{0})_{f_{0}},\mathbb{R}^{n})$ and
$|D\accentset{\leftarrow}{f}_{0}^{-1}|\leq\frac{C(f_{0})}{H_{0}\circ\accentset{\leftarrow}{f}_{0}^{-1}}.$
By Proposition 6.7 we have that $\accentset{\leftarrow}{f}_{0}^{-1}:(U_{0})_{f_{0}}\to F^{-1}B_{0}$ is analytically quasiconformal. Thus $\accentset{\leftarrow}{f}\circ\accentset{\leftarrow}{f}_{0}^{-1}:(U_{0})_{f_{0}}\to B_{0}$ and $P_{0}:=(\accentset{\leftarrow}{f}\circ\accentset{\leftarrow}{f}_{0}^{-1})^{-1}:B_{0}\to(U_{0})_{f_{0}}$ are analytically quasiconformal (again by Proposition 6.7, since $P_{0}$ is 1-Lipschitz). Thus $\accentset{\leftarrow}{f}^{-1}|_{B_{0}}=\accentset{\leftarrow}{f}_{0}^{-1}\circ P_{0}$ is analytically quasiconformal and moreover
$|D\accentset{\leftarrow}{f}^{-1}|$ $=|D(\accentset{\leftarrow}{f}_{0}^{-1}\circ P_{0})|\leq|D\accentset{\leftarrow}{f}_{0}^{-1}|\circ P_{0}|DP_{0}|$
$\leq\frac{C(f_{0})}{H_{0}\circ\accentset{\leftarrow}{f}^{-1}}\cdot C(f)\frac{H_{0}\circ\accentset{\leftarrow}{f}^{-1}}{C(f_{0})H\circ\accentset{\leftarrow}{f}^{-1}}=\frac{C(f)}{H\circ\accentset{\leftarrow}{f}^{-1}}\quad\textrm{on }B_{0}.$
By repeating this argument for all points $z_{0}\in\Omega_{f}\setminus A$ it follows that $\accentset{\leftarrow}{f}^{-1}|_{\Omega_{f}\setminus A}\in N_{loc}^{1,n}(\Omega_{f}\setminus A,\mathbb{R}^{n})$ with
$|D\accentset{\leftarrow}{f}^{-1}|\leq\frac{C(f)}{H\circ\accentset{\leftarrow}{f}^{-1}}\quad\textrm{a.e. on }\Omega_{f}\setminus A.$ (6.5)
Fix a smooth function $\varphi:\mathbb{R}\to[0,1]$ such that $\varphi=1$ in a neighbourhood of 0, $\operatorname{spt}\varphi\subset(-1,1)$ and $\operatorname{LIP}(\varphi)\leq 2$. For each $\varepsilon>0$ set
$G_{\varepsilon}:=\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)f\circ b+\Big(1-\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)\Big)\accentset{\leftarrow}{f}^{-1}.$
Note that $G_{\varepsilon}=f\circ b$ in a neighbourhood of $A$, $G_{\varepsilon}=\accentset{\leftarrow}{f}^{-1}$ in $\Omega_{f}\setminus B(A,\varepsilon)$, and $G_{\varepsilon}\in N_{loc}^{1,n}(\Omega_{f},\mathbb{R}^{n})$. [^3] Indeed, for all points in $B(A,\varepsilon)\setminus A$, $G_{\varepsilon}$ is a composition of maps in $W^{1,n}$.
Writing $G_{\varepsilon}=\accentset{\leftarrow}{f}^{-1}+\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)(f\circ b-\accentset{\leftarrow}{f}^{-1})$ we have
$|DG_{\varepsilon}|\leq$ $|D\accentset{\leftarrow}{f}^{-1}|+|f\circ b-\accentset{\leftarrow}{f}^{-1}|\Big|D\varphi\circ\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)\Big|$
$+\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)|D(f\circ b-\accentset{\leftarrow}{f}^{-1})|$
$\leq$ $|D\accentset{\leftarrow}{f}^{-1}|+|f\circ b-\accentset{\leftarrow}{f}^{-1}|\cdot 2\chi_{(-1,1)}\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)\frac{|D(f\circ b-\accentset{\leftarrow}{f}^{-1})|}{\varepsilon}$
$+\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)|D(f\circ b-\accentset{\leftarrow}{f}^{-1})|$
$\leq$ $|D\accentset{\leftarrow}{f}^{-1}|+3|D(f\circ b-\accentset{\leftarrow}{f}^{-1})|.$
Moreover
$G_{\varepsilon}-\accentset{\leftarrow}{f}^{-1}=\varphi\Big(\frac{f\circ b-\accentset{\leftarrow}{f}^{-1}}{\varepsilon}\Big)(f\circ b-\accentset{\leftarrow}{f}^{-1})$
which converges to 0 locally uniformly. Together with the uniform bound $|DG_{\varepsilon}|\leq|D\accentset{\leftarrow}{f}^{-1}|+3|D(f\circ b-\accentset{\leftarrow}{f}^{-1})|$ (independent of $\varepsilon$), this implies that $\accentset{\leftarrow}{f}^{-1}\in N_{loc}^{1,n}(\Omega_{f},\mathbb{R}^{n})$. Since $\mathcal{H}^{n}(A)=0$, (6.5) implies the bound on of the minimal upper gradient in the claim. The inequality $|D\accentset{\leftarrow}{f}^{-1}|^{n}\leq K_{I}K_{O}\mathbf{J}(\accentset{\leftarrow}{f}^{-1})$ follows from Proposition 6.7. This completes the induction step and the proof. ∎
We close the section with the proof of Theorem 1.4.
Proof of Theorem 1.4.
The first claim (1) is contained in Lemma 6.1, Proposition 6.2 and Proposition 6.4. Indeed, $n$-rectifiability follows from the fact that $\Omega_{f}$ is the image of a Sobolev map from an open set in $\mathbb{R}^{n}$ with property $(N)$, upper Ahlfors $n$-regularity proved in Proposition 6.2, and the infinitesimal Poincaré inequality in Proposition 6.4.
The second claim (2) follows by combining Proposition 6.6 and Theorem 6.8. Indeed, the analytic quasiconformality of the inverse $\accentset{\leftarrow}{f}^{-1}$, given by Theorem 6.8, is equivalent with the modulus inequality
$\operatorname{Mod}_{n}\accentset{\leftarrow}{f}(\Gamma)\leq K_{I}K_{O}\operatorname{Mod}_{n}\Gamma$
for path families in $f(\Omega)$. ∎
Appendix A Maximal Hilbertian factor of Almgren space
Throughout this section we fix integers $d,n\geq 1$. By convention $\mathcal{A}_{0}(\mathbb{R}^{n})=\{0\}$. Recall the barycenter map $b:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathbb{R}^{n}$ (1.6), and define
$\mathcal{Z}_{d}:=\{\llbracket\bar{x}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n}):\ b(\llbracket\bar{x}\rrbracket)=0\}.$
We let $b_{d}:(\mathbb{R}^{n})^{d}\to\mathbb{R}^{n}$ be the linear map $\bar{x}\mapsto\frac{x_{1}+\cdots+x_{d}}{d}$ and note that $b(\llbracket\bar{x}\rrbracket)=b_{d}(\bar{x})$. Given $\llbracket\bar{x}\rrbracket=\llbracket x_{1},\ldots,x_{d}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n})$, set $\llbracket\bar{x}-b(\bar{x})\rrbracket:=\llbracket x_{1}-b(\bar{x}),\ldots,x_{d}-b(\bar{x})\rrbracket$ and note that
$\llbracket\bar{x}-b(\bar{x})\rrbracket\in\mathcal{Z}_{d}\quad\textrm{for all }\llbracket\bar{x}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n}).$
Denote $\pi_{\mathcal{Z}}:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathcal{Z}_{d}$ the map $\llbracket\bar{x}\rrbracket\mapsto\llbracket\bar{x}-b(\bar{x})\rrbracket$.
The factorization in Proposition A.1 below reflects the fact that the diagonal of $(\mathbb{R}^{n})^{d}$ is the invariant set of the action of $S_{d}$. Indeed, it is easy to see that $\mathcal{L}$ maps $\operatorname{diag}\mathcal{A}_{d}(\mathbb{R}^{n})$ to the first factor.
Proposition A.1.
Equip $\mathbb{R}^{n}\times\mathcal{Z}_{d}$ with the Euclidean product metric. Then the map $\mathcal{L}:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathbb{R}^{n}\times\mathcal{Z}_{d}$ given by
$\mathcal{L}(\llbracket\bar{x}\rrbracket)=(\sqrt{d}b(\llbracket\bar{x}\rrbracket),\llbracket\bar{x}-b(\llbracket\bar{x}\rrbracket)\rrbracket)$
is an isometric homeomorphism. Moreover
  • (i)
    $\mathcal{Z}_{d}$ does not contain any geodesic line (i.e. isometry $\mathbb{R}\to\mathcal{Z}_{d}$).
  • (ii)
    there exists an open map $F:\mathcal{A}_{d-1}(\mathbb{R}^{n})\to\mathcal{Z}_{d}$ such that $F^{-1}(p)\leq d$ for all $p\in\mathcal{Z}_{d}$, and $\ell(F\circ\gamma)=\ell(\gamma)$ for all curves $\gamma$ in $\mathcal{A}_{d-1}(\mathbb{R}^{n})$.
We remark that (i) and (ii) are not needed anywhere in this paper, but are included out of independent interest. Property (i) states that $\mathbb{R}^{n}$ is the maximal Hilbert space factor of $\mathcal{A}_{d}(\mathbb{R}^{n})$, while property (ii) says that $\mathcal{Z}_{d}$ admits a branched covering from $\mathcal{A}_{d-1}(\mathbb{R}^{n})$.
Before providing the proof we record the following immediate corollary, which is used in Section 5.
Corollary A.2.
  • (i)
    The barycenter map (1.6) is $1/\sqrt{d}$-Lipschitz;
  • (ii)
    If $f:U\to\mathcal{A}_{d}(\mathbb{R}^{n})$ is approximately differentiable at $x$, then we have the inequalities
    $\|D_{x}f\|\leq\sqrt{d\|D_{x}(b\circ f)\|^{2}+\|D_{x}(\pi_{\mathcal{Z}}\circ f)\|^{2}},$
    $\max\{\sqrt{d}\|D_{x}(b\circ f)\|,\|D_{x}(\pi_{\mathcal{Z}}\circ f)\|\}\leq\|Df\|.$
Proof.
For each $\llbracket\bar{x}\rrbracket,\ \llbracket\bar{y}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n})$ we have
$d|b(\llbracket\bar{y}\rrbracket)-b(\llbracket\bar{x}\rrbracket)|^{2}\leq d|b(\llbracket\bar{y}\rrbracket)-b(\llbracket\bar{x}\rrbracket)|^{2}+d^{2}_{\mathcal{A}}(\llbracket\bar{x}-b(\bar{x})\rrbracket,\llbracket\bar{y}b(\bar{y})\rrbracket)=d_{\mathcal{A}}^{2}(\llbracket\bar{y}\rrbracket,\llbracket\bar{x}\rrbracket).$
This proves (i). To see (ii), simply apply the isometric splitting of Proposition A.1 to the observation that the operator norm of a linear map $L:\mathbb{R}^{m}\to\mathbb{R}^{a}\times\mathbb{R}^{b}$ satisfies
$\max\{\|P_{a}\circ L\|,\|P_{b}\circ L\|\}\leq\|L\|\quad\textrm{and}\quad\|L\|\leq\sqrt{\|P_{a}\circ L\|^{2}+\|P_{b}\circ L\|^{2}}$
Proof of Proposition A.1.
Let $\llbracket\bar{x}\rrbracket=\llbracket x_{1},\ldots,x_{d}\rrbracket,\llbracket\bar{y}\rrbracket=\llbracket y_{1},\ldots,y_{d}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n})$. By relabeling we may assume that
$d_{\mathcal{A}}^{2}(\llbracket\bar{x}\rrbracket,\llbracket\bar{y}\rrbracket)^{2}=\sum_{j}^{d}|x_{j}-y_{j}|^{2}.$
Then also
$d_{\mathcal{A}}(\llbracket x-b(\bar{x})\rrbracket,\llbracket\bar{y}-b(\bar{y})\rrbracket)^{2}=\sum_{j}^{d}|x_{j}-y_{j}-b(\bar{x}-\bar{y})|^{2}.$ (A.1)
Indeed, if $\sum_{j}^{d}|x_{j}-y_{\sigma(j)}-b(\bar{x}-\bar{y})|^{2}<\sum_{j}^{d}|x_{j}-y_{j}-b(\bar{x}-\bar{y})|^{2}$ for some $\sigma\in S_{d}\setminus\{id\}$, expanding the squares and eliminating the terms $|b(\bar{x}-\bar{y})|^{2}$ from both sides we obtain:
$\sum_{j}^{d}[|x_{j}-y_{\sigma(j)}|^{2}-2\langle x_{j}-y_{\sigma(j)},b(\bar{x}-\bar{y})\rangle]$ $<\sum_{j}^{d}[|x_{j}-y_{j}|^{2}-2\langle x_{j}-y_{j},b(\bar{x}-\bar{y})\rangle]$
$\Longleftrightarrow$
$\sum_{j}^{d}|x_{j}-y_{\sigma(j)}|^{2}$ $<\sum_{j}^{d}[|x_{j}-y_{j}|^{2},$
since
$\sum_{j}^{d}\langle x_{j}-y_{\sigma(j)},b(\bar{x}-\bar{y})\rangle=\sum_{j}^{d}\langle x_{j}-y_{j},b(\bar{x}-\bar{y})\rangle.$
This would be a contradiction, and thus (A.1) is established.
Denoting $z=(x_{1}-y_{1},\ldots,x_{d}-y_{d})\in(\mathbb{R}^{n})^{d}$, we recall that the average satisfies
$\sum_{j}^{d}\|z_{j}-b(z)\|^{2}=\sum_{j}^{d}\|z_{j}\|^{2}-d\|b(z)\|^{2},$
which yields
$d_{\mathcal{A}}(\llbracket x-b(\bar{x})\rrbracket,\llbracket\bar{y}-b(\bar{y})\rrbracket)^{2}=d_{\mathcal{A}}^{2}(\llbracket\bar{x}\rrbracket,\llbracket\bar{y}\rrbracket)^{2}-d\|b(\bar{x})-b(\bar{y})\|^{2}.$
This proves that $\mathcal{L}$ is an isometry. Since it is clearly a surjection, it is an isometric homeomorphism. Properties (i) and (ii) in the claim are proved in Lemmas A.3 and A.4, respectively. ∎
Lemma A.3.
There does not exist an isometric embedding $\mathbb{R}\to\mathcal{Z}_{d}$.
In the proof we use the projection map $\pi:(\mathbb{R}^{n})^{d}\to\mathcal{A}_{d}(\mathbb{R}^{n})$ which is 1-BLD, i.e. length preserving, discrete and open (cf. [19] for the terminology). Note in particular that $\pi|_{\ker(b_{d})}:\ker(b_{d})\to\mathcal{Z}_{d}$ is also 1-BLD.
Proof.
Suppose $\gamma:\mathbb{R}\to\mathcal{Z}_{d}$ is an isometry. Since $\pi|_{\mathcal{Z}_{d}}$ is 1-BLD, by [19, Lemma 2.8] there exist an affine geodesic line $\tilde{\gamma}(t)=a+tv$ in $\ker(b_{d})\subset(\mathbb{R}^{n})^{d}$ so that $\pi\circ\tilde{\gamma}=\gamma$. It follows that $a,v\in\ker(b_{d})$, and $|v|=1$. Given $t\in\mathbb{R}$ and $\sigma\in S_{d}$ we have
$(2t)^{2}=d_{\mathcal{A}}(\gamma_{-t},\gamma_{t})^{2}\leq\sum_{j}^{d}|\tilde{\gamma}_{j}(-t)-\tilde{\gamma}_{\sigma(j)}(t)|^{2}=\sum_{j}^{d}|t(v_{j}+v_{\sigma(j)})-(a_{j}-a_{\sigma(j)})|^{2}.$
Diving by $t^{2}$ and taking the limit $t\to+\infty$ we obtain
$4\leq\sum_{j}^{d}|v_{j}+v_{\sigma(j)}|^{2}=\sum_{j}^{d}[|v_{j}|^{2}+2\langle v_{j},v_{\sigma(j)}\rangle+|v_{\sigma(j)}|^{2}]=2+2\sum_{j}^{d}\langle v_{j},v_{\sigma(j)}\rangle,$
or
$1\leq\sum_{j}^{d}\langle v_{j},v_{\sigma(j)}\rangle,$
which can only hold if $v_{j}=v_{\sigma(j)}$ for $j=1,\ldots,d$. Since $\sigma\in S_{d}$ is arbitrary this implies that $v\in\operatorname{diag}(\mathbb{R}^{n})^{d}$, but since $v\in\ker(b_{d})$ we obtain $v=0$, a contradiction. This completes the proof. ∎
Lemma A.4.
There exists $\lambda\in(0,1)$ such that the map $F:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathcal{Z}_{d+1}$ given by
$F(\llbracket\bar{x}\rrbracket)=\llbracket x_{1}-\lambda b(\bar{x}),\ldots,x_{d}-\lambda b(\bar{x}),d(\lambda-1)b(\bar{x})\rrbracket,\quad\llbracket\bar{x}\rrbracket=\llbracket x_{1},\ldots,x_{d}\rrbracket,$
is open, $F^{-1}(p)\leq d+1$ for all $p\in\mathcal{Z}_{d+1}$, and $\ell(F\circ\gamma)=\ell(\gamma)$ for all curves $\gamma$ in $\mathcal{A}_{d}(\mathbb{R}^{n})$. Moreover, $F|_{\mathcal{Z}_{d}}$ is an isometric embedding.
Proof.
Given $\lambda\in[0,1]$, consider the linear map $A_{\lambda}:(\mathbb{R}^{n})^{d}\to(\mathbb{R}^{n})^{d+1}$ given by
$A_{\lambda}(\bar{x})=(x_{1}-\lambda b(\bar{x}),\ldots,x_{d}-\lambda b(\bar{x}),d(\lambda-1)b(\bar{x})),\quad\bar{x}=(x_{1},\ldots,x_{d}).$
Note that $b_{d+1}\circ A_{\lambda}=0$, which implies that $\operatorname{Im}(A_{\lambda})\subset\ker(b_{d+1})$. Indeed,
$(d+1)b_{d+1}(A_{\lambda}(\bar{x}))$ $=d(\lambda-1)b_{d}(\bar{x})+\sum_{j}^{d}(x_{j}-\lambda b_{d}(\bar{x}))$
$=d(\lambda-1)b_{d}(\bar{x})+db_{d}(\bar{x})-\lambda db_{d}(\bar{x})=0.$
Moreover we may calculate
$|A_{\lambda}(\bar{x})|^{2}$ $=d^{2}(\lambda-1)^{2}|b_{d}(\bar{x})|^{2}+\sum_{j}^{d}|x_{j}-\lambda b_{d}(\bar{x})|^{2}$
$=d^{2}(\lambda-1)^{2}|b_{d}(\bar{x})|^{2}+\sum_{j}^{d}[|x_{j}|^{2}-2\lambda\langle x_{j},b_{d}(\bar{x})\rangle+\lambda^{2}|b_{d}(\bar{x})|^{2}]$
$=|\bar{x}|^{2}+|b_{d}(\bar{x})|^{2}(d^{2}(\lambda-1)^{2}-2\lambda d+\lambda^{2}d).$
Denoting $P(\lambda)=d^{2}(\lambda-1)^{2}-2\lambda d+\lambda^{2}d$ we have $P(0)=d^{2}$ and P(1)=dP(1)=-d. Since $P$ is continuous, it takes the value 0 for some $\lambda\in(0,1)$. We fix this $\lambda$ for the rest of the proof and drop it from the subscript. Thus $A:(\mathbb{R}^{n})^{d}\to(\mathbb{R}^{n})^{d+1}$ is an isometric linear map with image $\ker(b_{d+1})$.
Consequently the map $F:\mathcal{A}_{d}(\mathbb{R}^{n})\to\mathcal{Z}_{d+1}$ given by
$F(\llbracket\bar{x}\rrbracket)=\llbracket A(\bar{x})\rrbracket$
is well-defined and satisfies $\ell(F\circ\gamma)=\ell(\gamma)$ for all curves $\gamma$ in $\mathcal{A}_{d}(\mathbb{R}^{n})$. Openness follows from the definition of quotient topologies, while for $\llbracket\bar{x}\rrbracket\in\mathcal{Z}_{d}$ we have $F(\llbracket\bar{x}\rrbracket)=\llbracket x_{1},\ldots,x_{d},0\rrbracket$ which is easily seen to be an isometry.
It remains to prove that $F^{-1}(p)\leq d+1$ for all $p$. If $p=F(\llbracket\bar{x}\rrbracket)$, then $b_{d}(\bar{x})$ can take at most $d+1$ distinct values (corresponding to the unordered components of $p$). But if $F(\llbracket\bar{x}\rrbracket)=F(\llbracket\bar{y}\rrbracket)$ and $b_{d}(\bar{x})=b_{d}(\bar{y})$, then $\bar{x}=\sigma\bar{y}$ for some $\sigma\in S_{d}$. This means that there can be at most $d+1$ distinct points $\llbracket\bar{x}\rrbracket\in\mathcal{A}_{d}(\mathbb{R}^{n})$ so that $F(\llbracket\bar{x}\rrbracket)=p$. ∎
Remark A.5.
The proof shows that in fact $F^{-1}(p)$ is the number of distinct points in $p=\llbracket p_{1},\ldots,p_{d+1}\rrbracket$.